A Complete Bibliography of Publications by, and about, Frank W. J. Olver

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

04 November 2023 Version 1.00

Abstract

This bibliography records publications of Frank W. J. Olver.

Title word cross-reference

(2 + 1) [FBZ21]. 1 [GS90]. C_{ν}(x) [EL98]. d^{4}u/dz^{4} + [z^{2}(d^{2}u/dz^{2}) + az(du/dz) + bu] = 0 [Par75]. Δu − u = f [Sie90]. Γ [Duv83]. H_{v}^{(1)} [Dai94]. kp = a^{2} + Db^{2} [GM60]. R^{3} [CV14]. μ [Par74]. n [Kam90]. P_{n}^{−m}(cosh z) [Fre90b, SW88]. Q_{n}^{−m}(cosh z) [Fre90a]. t [IK76]. u^{iv} + λ^{2}[(z^{2} + c)u′′ + azu′ + bu] = 0 [PW80]. |z| [PW80].

-dimensional [FBZ21]. -ième [Par74]. -th [Par74]. -transform [Dai94]. -variable [Kam90].

135 [Sch69]. 1964 [Wil64]. 1966 [Olv67a]. 1982 [Olv84]. 1984 [Olv86a]. 1987 [Tur89]. 1988 [Olv89a]. 1989 [Olv91a]. 1999 [DIW00].

2-Peakon [CYH22]. 21st [Ano11a].

34 [Fra81].

5-cyclidic [CV14]. 50th [Gau94].

60 [BHO^{+}67]. 65th [Won90].

75-20 [Boe76].

81f [Fra81]. 838 [Fab04]. 877 [Kod08]. 8th [IS87].

A. [Olv65a]. Abschätzung [Ber77]. Abstract [Olv54b]. Abundant [SQHY08, SHY11]. Academic [Olv89a]. Accelerating [BSD19, CC90]. Ackerberg [Olv78d]. across [PM81]. Addendum [Fra81]. Advances [Olv98a]. aerodynamic [Olv49]. Age [BCLO11]. aid [Ask90]. Airy [FO99, Fab04, FLO04b, FLO04a, GS14, HB75, Koh79, LO93, Olv10, Smi90, VRZG96]. algebra [KM14]. Algebraic [ROAW10, Olv90b]. Algebro [YX14]. Algebro-geometric [YX14]. ALGOL [BHO^{+}67]. Algorithm [FLO04b, Zah90, Cas80, CO80, Olv64b, Olv67c, SL72, Van79, Fab04, Kod08]. Algorithms [OS72, CO86, Olv80b, Olv85, Olv84]. Amsterdam [Olv67a]. analogue [Duv83]. Analysis [CO51, Olv84, Tur89, Won90, Olv64b, Olv65c, Olv65b, Olv75b, Olv82, Olv83b, Olv04, Zan95, Olv63b]. analytic [ROAW10]. Anniversary [Gau94]. Anticipated [Ano08]. Anwendungsgebiete [Sch69]. Ap [Olv82]. application [BD86, FW80b, Olv64d, Olv65e, Olv00, Olv65c]. Applications [Ano11a, MV90, Won83, Wil64, Zan95, Olv89a]. Applied [MW92, Dun14, Mil52, Zay11]. approach [CC85, Olv78c, Smi90]. approximate [Hea72]. Approximating [SV90a]. approximation [BG88, IKS90, MW83, Olv61a, Olv68, Sik83]. Approximations [Olv63a, SV94, WH86a, Bal79, LF90, LO78, Olv74c, Olv78b, Olv80a, Olv80c, OW10, Smi86, Olv72]. Arbitrary [OS65, Hea75, Olv77a]. Argument [Kod08, Olv64d]. arising [Iga85]. Arithmetic [Hay09, IS87, Olv87, OT87, Olv09, Swa90, CO87, COT89, CT89, LO90, Olv78c, Olv83b, Olv83c, Olv86b, Olv90b, SOL81]. Army [Wil64]. asimptoticheskie [Olv78a]. Asimptotika [Olv90a]. Askey [AGIN94, LW14]. aspects [DR90]. Associated [OS83, Boy90, Olv60, SS90]. Asymptotic [BE92, Boy90, Dun90, Fre90a, GO77, Jon93, LO78, OCK^{+}95, Olv64c, OS65, Olv74a, Olv80a, Olv83a, Olv94a, Olv97a, OW10, RCB90, Sie90, SFB90, Tem92, Wil64, Won90, BG88, BE84, BD86, Con90, Dai94, Dun89, FO99, FLS14, FW85, FW88, Fre90b, FW80b, GS90, HC91, Jon90, KD14, LR86, Mil52, OO94, OO98, Olv51, Olv52b, Olv53, Olv54a, Olv54b, Olv54c, Olv56a, Olv56b, Olv58, Olv59b, Olv64d, Olv65e, Olv65d, Olv75b, Olv75d, Olv76b, Olv78a, Olv91b, Olv91c, Olv93a, Olv93b, Olv95, Olv98a, Olv99, Par75, PW80, Rei90, Sal83, Sal84, Sha88, SW88, Ski93, Son80, Tem78, Tem82, Tem83, Tem89, Tem90, Urs87, Wim80, Won81, Won83, Won89, Olv67a]. asymptotically [Bal79]. Asymptotics [Ber90, CI91, Lud83, MW92, Mey80, Olv74b, Olv90a, Won14a, Boe91, CF90, FLO04a, Han90, LW14, Nem14, Olv70a, Par90, Wim90, Wim75]. asymptotique [Duv83]. asymptotischen [Ber77]. August [Gau94]. Automatic [CO51, Sch69]. Award [Ano11c]. axial [Wan90].

Bäcklund [SHY11]. Backward [OS72]. Band [Sch69]. barrier [Hea72]. Barriers [Hea75, Dam83, Dam86]. Bateman [Olv46]. Bauer [Sch69]. Beams [BEZHN15, BSD19, WYX^{+}23]. behavior [FBZ21]. behaviour [GO77, Par75]. Berücksichtigung [Sch69]. besonderer [Sch69]. Bessel [F.63, BS14, Bha73, Boe91, Bot77, Cha71, Dai94, DL91, GS14, GO77, IK76, Ism77, Kno70, KT60, LMS73, LO93, McC77, Mul79, Nem14, OR71, Olv50, Olv51, Olv52b, Olv54a, Olv60, Olv62, Olv64a, OM10, SP73, Sal83, Sal84, Sta83, SFB90, Wim90, YK60]. Bessel-function [Boe91]. beta [FW80b]. beta-type [FW80b]. Beyer [MW92]. Beyond [CO84]. bifurcation [AHS^{+}18]. binomial [GN14]. binomial-type [GN14]. Birkhäuser [Olv84]. Birthday [Won90]. blow [LOQZ14]. blow-up [LOQZ14]. Bogoyavlenskii [Alj21]. Boisvert [Cro11, Tho11]. Book [Cro11, F.63, Jor80a, Olv65a, Olv67a, Olv72, Olv84, Olv86a, Olv86e, Olv89a, Olv91a, Olv97b, Sch69, Wim75]. Boston [Olv84, Olv86a]. both [Dun91, Olv75a, Olv80c]. bound [Dun14, EL98]. Boundary [Lak72, WH86a, WH86b, AO90]. bounded [Hea75]. Bounds [Fre90a, Olv63a, Olv64c, OS65, Olv67b, OW82, Olv88, BG88, Fre90b, McC77, MW83, Olv61a, Olv64d, Olv65d, Olv68, Olv74c, Olv76a, Olv80a, Olv83c, Olv86b, Olv86d, SW88]. breaking [Alj21]. bridge [KW90]. British [Gau94]. Brown [GN14].

calculation [AB87, OO95b]. calculations [Olv49]. Camassa [LOQZ14]. Cambridge [Olv65a]. case [Boy90]. Cauchy [Olv14b]. Center [Wil64]. Centrum [Olv67a]. Century [Ano11a, Lid01, Gau94]. certain [LLS80, Mul73, Olv49, SL72]. Chair [MW92]. Change [Olv14b]. Chapman [UO61]. Chapter [BE84]. Charles [Cro11, Tho11, Olv04]. Chebyshev [CC90, SL72]. Chester [SS90]. Chronicle [Lid01]. circle [WYX^{+}23]. Clark [Cro11, Tho11]. class [BD86, Hin79, LLS80, Tem82]. Classical [Olv97b]. Clenshaw [Olv04]. Closed [Olv87, Olv09]. Closure [LO90, Olv89b]. Coalescing [Dun90, Con90, LR86]. coefficient [Boy90, Dam83]. coefficients [FW80a, Fra81, Olv94a, Wim90]. coincident [Bal79]. Collection [LO70]. Columbia [Gau94]. combinatorial [Mac75]. Como [IS87]. Comp [Fra81]. Comparison [KYD20]. complete [Dam86, Fre90b, Ism77]. Complex [Dun90, Kod08, FLO04a, GS14, Olv65d, Olv78b, Olv83b, Olv86b, Par90, YX14]. Computation [FLO04a, Ral65, CH90, DR90, Gau94, Sch69, Tem89, Olv86a]. Computational [CF90, Won90, Gau94]. Computer [IS87, Olv84, Olv87, Olv09, Swa90]. computers [Olv83c]. Computing [CO51, CGM^{+}61, FGM^{+}57, FGM^{+}70, SKL93, VRZG96]. conclusions [Olv89c]. condition [Dam86]. conditions [Olv78d]. Conducted [Wil64]. Conference [Ano11a, SKL93, Won90]. confluent [Dun89, Gat90, Olv91b, Olv93b, Tem78]. Conical [Dun91]. conjecture [GN14, QW88]. Conjectured [Jor80a, Jor80b]. Connection [Olv77a, Olv77b, Olv78b, Olv80d, PM81]. connection-formula [Olv80d]. conservation [AHS^{+}18]. considerations [Hea72]. constants [QW88]. containing [Bal79, Olv56a, Sik83]. Continental [Dav90]. Continuous [SV94]. contractions [KM14]. control [WYX^{+}23]. convergence [CC90]. Convergent [DL91, FLS14]. Converging [Olv90c, Olv92, Mil52]. convexity [BS14]. convolution [Sik83, Wim90]. Corrigendum [Olv79, Fra81]. Coulomb [Olv65a]. count [Hay09]. cross [Mul79, SP73]. cross-product [Mul79, SP73]. cumulative [Tem82]. Curtis [Olv65a]. cut [OS83]. cyclidic [CV14]. cylinder [EL98, LO78, Mil52, Olv59b, Olv61b, Olv64d, Olv80c]. Cylindrical [Kod08, KW90].

D [Olv84, Olv89a]. Daniel [Cro11, Tho11]. data [Shi90]. DC [LO70]. decaying [Olv14b]. Dedication [AGIN94]. definite [Olv68, SL72]. degenerate [LF90]. degree [Olv52a]. derivation [Cha71]. derivatives [Bus76, GS14]. descent [Kam90]. descents [MV90, Olv70b, Olv70c]. Description [BHO^{+}67]. Determination [SP73, Mil52]. developments [Olv82]. deviations [FW80b]. Difference [Olv89a, Zah90, Cas80, DV14, HT94, Olv67b, Olv67d, Olv00, Van79]. different [Mul79]. differentiable [Wal91]. Differential [CO51, Dun90, FH90, OS65, Bal79, BD86, Boy90, Dav78, DV14, Dun14, FLO04a, FLS14, Hin79, OO94, OO95b, OO95a, OO98, Olv50, Olv54c, Olv56a, Olv56b, Olv58, Olv59a, Olv65e, Olv75b, Olv75c, Olv76a, Olv77a, Olv77b, Olv77c, Olv80d, Olv93a, Olv93b, Olv94a, Olv97a, Olv98a, Olv98b, Olv99, Par75, PW80, Par92, SV90b, Wil64]. differential-difference [DV14]. Digital [Ano11c, BCLO11, Ral65, Ano10a]. digits [Olv83c]. Dilcher [GN14]. dimensional [FBZ21]. dimensions [Mey90]. Diophantine [GM60]. discontinuous [Shi90]. Discrete [SV94, Sha88]. Discussions [CO51]. displaying [GM60]. distribution [DV14, GS90, HB75, IK76, Tem82]. Divergence [Col80]. Docev [Sal83, Sal84]. domain [Olv56a]. dont [Par74]. Double [Dun90]. Duffing [JTB^{+}13]. d’un [Par74, TL76]. d’une [Duv83].

economized [CO55]. edited [Cro11, Tho11]. edition [Olv84]. égal [Par74]. Eigenfunction [FH90]. eigenproblems [Boy81]. eigenvalues [Dav78, WP90]. Einzeldarstellungen [Sch69]. elastic [Wan90]. Elementary [RO10, Olv80b]. Elimination [OW82]. Elliptic [Car87, Car88, LF90]. end [Hay09]. endpoint [LR86]. Engineering [Olv89a]. Entwicklung [Ber77]. equal [Par74]. Equation [CYH22, KYD20, AHS^{+}18, Alj21, Bal79, CV14, Dav78, DV14, FLO04a, FBZ21, GM60, JTB^{+}13, LOQZ14, Par75, PW80, S.15, SQHY08, SHY11, Smi90, Waz07, Zay11, ZAAN16]. Equations [Bat46, CO51, Dun90, OS65, Olv89a, SV94, Alj21, BD86, Boy90, Cas80, Dun14, FLS14, HT94, Hin79, Iga85, LLS80, OO94, OO95b, OO95a, OO98, Olv50, Olv54c, Olv56a, Olv56b, Olv58, Olv59a, Olv65e, Olv67b, Olv67d, Olv75b, Olv75c, Olv76a, Olv77a, Olv77b, Olv77c, Olv80d, Olv93a, Olv93b, Olv94a, Olv97a, Olv98a, Olv98b, Olv99, Olv00, Par92, SV90b, Van79, Wil64]. equivalence [Koo14]. Errata [Sal84]. Erratum [Ski81]. Error [Fre90b, Fre90a, MW83, Olv61a, Olv63a, Olv64b, Olv64d, Olv64c, Olv65c, Olv65b, Olv65d, OS65, Olv68, Olv74c, OW82, Olv83b, Olv83c, Olv86b, Olv88, Pry84, Ral65, SW88, BG88, Dun14, Oli68, Olv75b, Olv76a, Olv78c, Olv80a, Olv82, Olv86d]. Errors [Olv56b, Olv90b]. est [Par74]. estimates [Tem90]. estimation [Oli68, Olv75d]. Étude [Duv83]. evaluate [SL72]. evaluating [SV90b]. Evaluation [Olv66, Con90, FW85, LO94a, LO94b, Olv50, Olv51, Olv52a, Olv86b]. Evolutionary [FBZ21]. Exact [AHS^{+}18, Alj21, KYD20, Hea72, S.15, SQHY08, SHY11]. Excellence [Lid01]. existence [Dav78, LF90]. exp [ZAAN16]. exp-function [ZAAN16]. Expansion [FH90, Fre90a, BE92, FO99, Fre90b, FW80b, HC91, LR86, Olv53, Olv54a, Olv54b, Olv95, PW80, SW88, Tem82, Wim80, Won81]. Expansions [CV14, DL91, OCK^{+}95, Olv64c, Ski80, BE84, Boy90, Dai94, Dun89, FLS14, FW88, KD14, Mil52, Olv51, Olv52b, Olv58, Olv59b, Olv64d, Olv65d, Olv76b, Olv91b, Olv91c, Olv94a, Sie90, Ski81, Ski93, Son80, SFB90, Tem78, Tem83, Tem89, Urs87, Won83, Won89, Olv67a]. explicit [SQHY08, SHY11]. Exponent [Dun90]. Exponential [Mey80, CO80, CLOT86, Han90, Olv91c, Olv94b, Wal91]. Exponentialintegrals [Ber77]. Exponentially [OO94, Olv93a, Olv93b, Olv91b, Olv91c, WP90]. Exponentially-improved [Olv93b]. exponentials [Olv89c]. Extended [SOL81, Koh79]. Extended-range [SOL81]. extension [Cas78, HT94, Oli68, Olv67c, Wan90].

F [Cro11, Sch69, Tho11, Wim75]. F. [F.63, Sch69, Won00]. Factorial [DL91]. Factors [Olv90c, Mil52, Olv92]. falling [Hay09]. Fermionization [YL12]. filter [IKS90]. finding [Iga85]. Finite [Wan90, Son80]. First [Olv63a, BS14, Koh79]. fitting [Olv89c]. flair [Olv04]. floating [CO84]. flows [KW90]. Fokas [CYH22]. fonction [Duv83]. Foolproof [Hay17]. formal [Sib90]. forms [GM60]. formula [Olv80d, Sha88]. formulae [Olv78b]. Formulas [AS64, Olv77a, Olv77b, Olv14b, Sta83]. FORTRAN [Olv91a]. Fourier [Son80, Won81]. Fourier-type [Won81]. fourth [CF90]. fourth-order [CF90]. Fractional [JTB^{+}13, Alj21, Olv77c, S.15, ZAAN16]. Frank [Cro11, Tho11, AGIN94, Ano13a, Ano13b, BB13, Ber14, Won90, Won13, Won14b]. French [Duv83, Par74]. Friedman [SS90]. Fritz [MW92]. Function [Fre90a, Olv91a, Boe91, CO80, EL98, FW80a, Fra81, Fre90b, Has60, Kno70, Koh79, LLS80, OR71, Olv91b, Olv93b, SP73, SW88, SV90a, Ste72, ZAAN16]. Functions [AS64, Ano11a, BCLO11, DL91, F.63, Fab04, FLO04b, Kod08, LO93, LO70, Olv62, Olv64a, Olv72, Olv74b, Olv97b, OLBC10, Tem92, AB87, Ask90, BS14, Bha73, BD86, Boy90, Bus76, CC85, Cha71, CLOT86, DIW00, Dun89, Dun91, FO99, FLO04a, Gat90, GS14, GO77, GS90, HB75, IKS90, Ism77, Jor80b, KD14, KT60, LMS73, LO78, LO94a, LO94b, McC77, Mil52, Mul73, Mul79, Nem14, Olv50, Olv51, Olv52b, Olv53, Olv54a, Olv54b, Olv59b, Olv60, Olv61b, Olv64d, Olv65e, Olv65d, Olv66, Olv75a, Olv75d, Olv78a, Olv80b, Olv80c, OS83, Olv85, Olv90a, Olv95, Olv00, Olv10, OM10, RO10, Sta83, SFB90, Tem78, Tem82, Tem89, VRZG96, Wal91, Wim80, Wim90, Won89, YK60, Zan95, Ano10a, Ano11c, Olv65a, Cro11, Jor80a, Tho11, Wim75, Ano11d, BL01]. fundamental [CV14, Koo14]. funktsii [Olv78a, Olv90a]. Further [Hea72, Olv82, Olv51].

Gallery [Ano10b, Ano11e]. Gamma [Tem92, AB87, FW80a, Fra81, Olv95]. Gauss [BSD19]. Gaussian [OW82]. Gegenbauer [Tem90]. General [Olv78b, Dai94, Olv80d, PM81, Par74, Par74, Olv65b]. generalised [Ojh73]. Generalization [Par94, Zah90]. Generalizations [Kno70]. Generalized [CLOT86, Olv89c, HB75, JTB^{+}13, OR71, Olv53, Olv91c, Olv94b, Sal83, Sal84, Wal91]. generating [Cha71, Olv80b]. Generation [BEZHN15, WYX^{+}23]. geometric [YX14]. Gevrey [Sib90]. given [Tur89]. Glasser [Boe76]. Global [LW14]. Graphs [AS64, Ask90]. Green [SV94, Olv61a, Olv78b, Smi86]. Greene [Olv84]. Grosswald [Sal83, Sal84]. Grundlehren [Sch69]. guard [Olv83c].

H [Olv67a, Olv84, Sch69]. H. [Olv46]. Hahn [Sha88]. half [Gau94]. half-century [Gau94]. Handbook [AS64, BCLO11, Cro11, OLBC10, Tho11, Sch69, Ano11d, BL01]. Hankel [FW85, Nem14]. harmonics [CV14]. Having [Dun90, BD86, Dun14, Olv65e, OS65, Olv77a, Sha88]. held [LO70, SKL93]. helicoidal [Wan90]. Her [F.63]. hierarchy [YX14]. High [FW80a, Fra81, GO77, Olv52a, Olv98b, Par92]. high-degree [Olv52a]. high-order [Olv98b, Par92]. High-precision [FW80a, Fra81]. Higher [CYH22, Hay09]. Higher-Order [CYH22]. Hilbert [Olv14b]. Holm [LOQZ14]. Home [Olv09]. Hong [DIW00]. Honor [Won90]. Honorees [Ano10b, Ano11e]. Householder [Sch69]. Hyperasymptotic [OO95a]. Hyperasymptotics [BH90]. hypergeometric [Dun89, Gat90, KD14, Olv53, Olv91b, Olv93b, Tem78].

identities [Koo14]. ième [Par74]. II [Olv72, Dam86, Mul73, OO94, Olv65c, Won00]. iii [F.63]. imaginary [WP90]. Implementation [OT87]. impliquant [TL76]. Improved [Olv76a, OO94, Olv91b, Olv91c, Olv93a, Olv93b]. impulse [IKS90]. Incomplete [Tem92, AB87]. Index [OT87, CO87, CT88, COT89, CT89, LO90, Olv86c, Olv90b]. inequalities [Bus76, Gat90, Olv61b]. Infinitely [Wal91]. infinity [Sie90]. Institute [Ano11b]. Integer [Olv64a]. integrable [LOQZ14]. Integral [Bat46, Ism77, Par94, SV94, Boe76, FW80b, OR71, Olv65c, Olv65b, Olv91b, Olv91c, Olv94b, Par90, Urs87, Won81]. intégrale [Duv83]. Integrals [Car87, Car88, Ski80, Urs90, Ber90, BE84, Bha73, Car90, Con90, GO77, LR86, Mul73, Olv68, SL72, Ski81, Ski93, Son80, SFB90, Tem82, Tem83, Wim80]. Integration [LO93]. interference [Olv49]. Interior [WH86a, Boy81, Shi90]. International [Ano11a, DIW00]. intrigued [Olv14a]. Introduction [BEZHN15, Olv97b, Olv78a]. introductory [COT89]. invariant [DR90]. Inversion [Tem92]. involve [Mul73]. involving [Bha73, SL72]. Irregular [OS65, Olv65e, Olv97a, OO94]. issue [AGIN94]. Italy [IS87]. IV [KD14].

J [AGIN94, Ano13b, BB13, Cro11, F.63, Sch69, Tho11, Wim75, Won90, Won00, Won13, Won14b]. Jacobi [BG88, CI91, Tem90]. Jet [Olv86a]. June [DIW00].

Kelvin [YK60]. kernels [Mul73]. Kind [Car87, Car88, BS14, HC91, Koh79]. kinks [Waz07]. Knuth [Olv84]. Kong [DIW00].

L [Boe76, Olv72, Sch69]. Laboratories [Olv91a]. Laguerre [FW88, Tem90]. Lakshmikantham [Olv89a]. Lancaster [Tur89]. Landen [Car90]. Laplace [Ski81, CV14, MW83, Olv68, Ski80, Son80, Tem83]. Large [F.63, Olv62, Dav78, Dun91, FLS14, FW80b, KD14, Nem14, Olv51, Olv52b, Olv54a, Olv54b, Olv54c, Olv58, Olv59a, Olv59b, Olv64d, Olv75a, Olv80c, PW80, Urs87, Urs90]. large-order [Nem14]. Lauwerier [Olv67a]. laws [AHS^{+}18]. layer [Shi90]. Lebesgue [QW88]. Lectures [Tur89]. Legendre [Jor80a, BD86, Bus76, Fre90b, Fre90a, Jor80b, Ojh73, Olv54b, Olv75a, OS83, Olv00, Par74, QW88, SW88, SOL81]. Level [CO87, COT89, OT87, CT88, CT89, LO90, Olv86c, Olv90b]. Level-Index [OT87, CO87, COT89, CT88, CT89, LO90, Olv86c, Olv90b]. Library [Ano10a, Ano11c, Olv91a]. lid [Dav90]. Liénard [AO90]. Limit [Boe76]. line [Boe91, Hay09]. Linear [Dun90, Olv59a, Olv88, Zah90, Bal79, BD86, Boy90, Cas78, Cas80, Dun14, HT94, OO95b, OO95a, OO98, Oli68, Olv54c, Olv56a, Olv56b, Olv58, Olv67b, Olv67d, Olv75b, Olv75c, Olv80d, Olv94a, Olv97a, Olv98a, Olv98b, Olv99, Shi90, SV90a, SV90b, Van79, Won14a, Olv86e]. Liouville [Boy81, Mul73, Olv61a, Olv78b, Smi86, SV94]. liquid [KW90]. Locating [VRZG96]. logarithmic [CLOT86, Wal91]. logarithms [Olv89c]. Lommel [Bha73, Ste72]. London [F.63]. Look [OT87]. Look-Up [OT87]. Lower [McC77]. Lozier [Cro11, Tho11]. Luke [Olv72].

M [Olv97b]. M. [Boe76]. Machinery [CO51]. Madison [Wil64]. magnetoconvection [RCB90]. Majesty [F.63]. Manchester [MW92]. manipulation [CC85]. Manna [GN14]. Marangoni [KW90]. March [SKL93]. mark [MW92]. Matched [OCK^{+}95, Ski93, WH86a]. Math [Ano11d, Fra81]. Mathematical [AS64, Ano10a, Ano11c, BL01, Cro11, Hay17, Olv65a, Olv97b, OLBC10, Tho11, CO55, Olv85, Olv91a]. Mathematician [Olv04, Ano13b]. Mathematics [Ano08, MW92, Olv89a, Olv14a, Gau94, Gau94, Wil64, Olv84]. Mathematisch [Olv67a]. mathematischen [Sch69]. maximum [Dam86, Urs90]. maximum-modulus [Urs90]. May [IS87, Wil64]. me [Olv14a]. Measure [Pry84]. Measurements [Lid01]. Meditations [Hay17]. meeting [LO70, MW92]. memoriam [Won14b]. Method [BEZHN15, Cas78, HT94, JTB^{+}13, MV90, Mil52, Olv50, Olv51, PM81, SV90b, Zay11, ZAAN16]. Methods [CGM^{+}61, FGM^{+}57, FGM^{+}70, Olv89a, Con90, Dun14, Kam90, Olv65c, Olv65b, Olv74a, Olv78a, Olv83a, ROAW10, RCB90, SV90a]. metody [Olv78a]. Microsoft [Olv91a]. Microsoft-FORTRAN [Olv91a]. Miller [Olv64b, Olv67c]. mod [GS90]. model [JTB^{+}13]. Moderate [F.63, Olv62]. Modern [CGM^{+}61, FGM^{+}57, FGM^{+}70]. modified [LOQZ14, Zay11, ZAAN16]. modifiée [Duv83]. modulus [Urs90]. monotonicity [Fre90b, Ism77, LMS73]. MR [Fra81]. Much [Ano08]. Much-Anticipated [Ano08]. multidimensional [MW83]. multiple [Olv77b, ZAAN16]. multiplicities [Olv77a]. multipliers [Hin79, OO95b].

national [LO70, Ano11b]. NBS [Lid01]. NBS/NIST [Lid01]. Necessary [Dam86]. Negative [Ver91]. Newton [Olv86d]. Nico [Olv97b]. NIST [Ano08, Ano11c, Ano11d, Lid01, OLBC10, Cro11, Tho11]. no [Fra81, Sha88]. Non [CYH22, Dav78]. Non-existence [Dav78]. Non-uniqueness [CYH22]. nondegenerate [LF90]. Nondiffracting [BEZHN15]. nonlinear [AHS^{+}18, Iga85, JTB^{+}13, LLS80, S.15, ZAAN16]. Nonlinearity [CYH22]. Nonnegative [Kod08]. Norfolk [SKL93]. normalized [BS14, SOL81]. Note [BSD19, Olv46, Olv53, OS72, Sal83, Cas80, FW85, LR86, Zay11, Sal84]. notebook [BE84]. Notes [CO51]. Novel [BEZHN15]. number [Hay09, Olv77a, Olv86c]. Numerical [AB87, CO51, KYD20, LO94a, LO94b, Olv63b, Olv67d, Olv89a, Tur89, Zan95, Cas78, CH90, LF90, OO98, Olv98a, Olv04, SV90b, Van79]. nye [Olv78a, Olv90a].

Obituaries [Won13]. Obituary

[Ano13a, UO61]. Observable [Mey90]. occurring [Boe91, Olv49]. odd [Hea75]. ODEs [LO93, SV90a]. off [Hay09]. Office [F.63]. OK [Col80]. Olmo [IS87]. Olver [CYH22, Cro11, F.63, Sch69, Tho11, Wim75, Won00, AHS^{+}18, Alj21, AGIN94, Ano13a, Ano13b, BB13, BEZHN15, Ber14, BSD19, Cas78, Cas80, Dun14, FBZ21, HT94, JTB^{+}13, KYD20, Oli68, S.15, SL72, SQHY08, SHY11, SV94, Van79, Ver91, WYX^{+}23, Waz07, Won90, Won13, Won14b, YL12, YX14, Zah90, Zay11, ZAAN16]. O’Malley [Olv78d]. one [Dun91, OO94, Olv56a, Olv65e]. Online [Ano08]. Operational [Cha71]. operations [CO87, Olv83c]. Operator [FH90]. operators [Ber90, CF90, Sik83, Ver91]. Order [CYH22, Dun90, FH90, Kod08, Olv64a, OS65, Bal79, BD86, Boy90, CF90, Cas80, Dav78, FLS14, GO77, HT94, Hea75, KM14, Nem14, OO95b, OO95a, Olv50, Olv51, Olv54a, Olv54b, Olv54c, Olv56a, Olv58, Olv59a, Olv65e, Olv67b, Olv67d, Olv75c, Olv76a, Olv77a, Olv77b, Olv77c, Olv80d, Olv98b, Par92, SV90a, SV90b, Van79].

Orders [Olv62, Mul79, Olv52b, Olv59b, F.63]. ordinary [Dun14, Hin79, OO94, OO98, Olv56b, Olv93a, Olv93b, Olv97a, Olv98a]. Orr [Rei90]. orthogonal [Sha88]. oscillating [Con90]. oscillations [KW90]. oscillatory [Olv14b]. Other [Hay17, Olv50, Olv91b, Zan95].

Package [Kod08]. page [Olv09]. paper [Olv46]. Papers [LO70, Won00]. parabolic [LO78, Mil52, Olv59b, Olv61b, Olv80c, Shi90]. paradox [Olv70a]. Parallel [LO93, SKL93, Tur89]. parameter [FLS14, Olv54c, Olv58, Olv59a, Sib90, Urs87, Urs90]. parameters [Dun91, KD14, Olv75a, Olv80c, Sik83]. Parsimonious [Lud83]. Part [Olv60, Won00, YK60, WP90]. Partial [OT87, Bus76, Dav90]. Partitions [GGM58, HC91]. Peakon [CYH22]. Pearcey [Par90, Par94]. penetration [Olv65c]. Period [KW90]. periodic [SQHY08]. perturbations [Olv76b]. perturbed [AO90]. phase [Olv65c, Olv65b, Olv74c, SV90a]. phase-integral [Olv65c, Olv65b]. Phenomenon [OCK^{+}95, Olv90c, Par92]. physics [SL72, Olv97b]. Pitman [Olv86a]. plane [AO90, Olv65d, Olv78b]. plasma [SL72]. Point [Dun90, Olv63a, Olv64c, Olv86e, WH86a, WH86b, BD86, CO84, Dun14, Han90, Kam90, Lak72, Olv56a]. points [Con90, Hea75, LR86, Olv65b, Olv75c, Olv76a, Olv77a, Olv77b, Olv77c, PM81]. Pole [Dun90, Boy81, Boy90]. polynôme [Par74, TL76]. Polynomial [Tem90, Olv86b, Par74]. Polynomials [CI91, BG88, Bot77, Cha71, CO55, DV14, DM14, FW88, GN14, IK76, LW14, Ojh73, Olv52a, SL72, Sal83, Sal84, Sha88, SOL81, SS90, Tem90]. Portrait [Ano10b, Ano11e]. Positive [LF90]. Posteriori [OW82].

Potential [Bat46, Dam83, Dam86]. powers [Par74, Ver91]. pp. [F.63, Olv65a]. Practical [Con90, Iga85, Olv98b]. precision [FW80a, Fra81, LO90, Olv89b]. Preface [TW14a, TW14b, TW14c]. Presented [LO70]. Press [Olv65a, Olv89a]. Preview [Ano08]. primes [GM60]. principle [Urs90]. Probabilistic [Sta83]. probabilities [FW80b]. Problem [Boe76, Boe91, Dam86, Hea72, Mac75, Olv80d, Rei90, Shi90]. Problems [Olv63a, Olv64c, AO90, Iga85, Lak72, MV90, Olv65c, Olv75d, Olv98b]. Proceedings [IS87, Wil64, MW92, DIW00, SKL93]. processes [Olv90b]. Processing [SKL93, Tur89]. product [Mul79, SP73]. products [Bha73, Wim90]. Professor [MW92]. Progress [Olv84]. proofs [Sta83]. Propagation [LO70, MV90]. properties [LMS73, Nem14, Wim90]. Property [Jor80a, Jor80b, Sib90]. pseudo [DM14]. pseudo-ultraspherical [DM14]. Publications [Lid01, Olv63b]. puissances [Par74, TL76].

Qiao [CYH22]. Quadratic [KM14, GM60]. quotients [Ism77].

R [Olv65a]. radius [BS14, KT60]. Raina [Jor80a]. Ramanujan [BE84, BE92]. range [SOL81]. Rank [OS65, OO94, Olv65e, Olv97a]. ratio [Olv95]. Rational [IKS90]. reaction [JTB^{+}13]. Receives [Ano11c]. reciprocals [CO86]. Recognizes [Ano11b]. Recurrence [CO51, OS72, Olv86a, Olv88, CC85, Cas78, Oli68, Olv64b]. recurrences [Won14a]. recursion [Ver91]. Reference [Ano08]. reformulation [Van79]. region [Bal79]. regular [BD86]. regularizing [Olv14b]. related [Dai94, FW80a, Fra81, LLS80, Olv10, Tem82]. Relations [CO51, Olv88, CC85, Cas78, Kno70, Oli68, Olv86a]. Relative [Pry84].

Releases [Ano08, Ano11d]. Reliable [CH90]. remainder [Sha88]. remainders [Jon90, Jon93]. removal [Dav90]. Representations [GM60, Ism77]. Research [Wil64]. Resonance [WH86b, Olv78d]. Restgliedes [Ber77]. results [Ojh73, Won83, Won89]. Resurgence [Olv00, Nem14]. retirement [MW92]. reversion [FO99]. Review [Cro11, F.63, Jor80a, Olv65a, Olv67a, Olv72, Olv84, Olv86a, Olv86e, Olv89a, Olv91a, Olv97b, Sch69, Wim75]. Riccati [Smi86, Smi90]. Richard [AGIN94]. Riemann [Has60]. rigid [Dav90]. Roger [UO61]. Ronald [Cro11, Tho11]. Root [CT89]. roots [CO86, Iga85]. Rosenau [CYH22]. Rounding [Olv90b]. Royal [Olv65a]. Rp [Olv82]. rule [AB87, Olv86d]. rules [Col80]. Russian [Olv78a, Olv90a]. Rutishauser [Sch69].

S [Fra81, Sch69]. saddle [Con90, Kam90, LR86]. Samelson [Sch69]. satisfying [DV14]. scale [Wim80]. scattering [Dav90]. School [Tur89]. Science [Olv84, Olv89a]. Scientific [SKL93]. Second [Car87, Dun90, FH90, OS65, Olv75c, Olv77c, Bal79, BE84, BD86, Boy90, Cas80, FLS14, HT94, KM14, OO95b, OO95a, Olv50, Olv54c, Olv56a, Olv58, Olv59a, Olv65e, Olv67b, Olv67d, Olv76a, Olv77a, Olv77b, Olv80d, Olv84, SV90a, SV90b, Van79]. Second-Order [Dun90, FH90, OS65, Olv75c, Olv77c, Bal79, BD86, Boy90, Cas80, FLS14, HT94, KM14, OO95a, Olv50, Olv58, Olv65e, Olv67b, Olv67d, Olv76a, Olv77a, Olv77b, SV90a, SV90b, Van79]. Selected [Lid01, Won00]. Self [BSD19]. Self-Accelerating [BSD19]. sequences [Wim90]. série [Par74, TL76]. Series [DL91, Ber90, BE92, Boe91, CC90, Jon93, Olv49, Olv94a, Par74, QW88, Sal83, Sal84]. several [Con90, Mey90]. Sharma [AHS^{+}18, Alj21, FBZ21, JTB^{+}13, KYD20, S.15, SQHY08, SHY11, Waz07, YL12, YX14, Zay11, ZAAN16]. shelf [Dav90]. shells [Wan90]. Short [Cro11]. SIAM [LO70, SKL93]. sign [Ste72]. simple [Boy90, Dun14, Olv65b, Zay11, ZAAN16]. Simplified [Olv86d]. Singh [Jor80a]. Singular [FH90, Han90, Mul73, WH86a, WH86b, Olv76b]. singularities [OO94]. Singularity [OS65, BD86, Olv65e, Olv97a]. Singularly [AO90]. Sixth [SKL93]. slowly [Olv14b]. small [Bal79, Urs87, WP90]. Smoothing [Par92]. Society [Olv65a]. solitary [SQHY08]. soliton [Alj21]. solitons [Waz07]. Solution [CO51, Cas78, Cas80, CV14, HT94, OO98, Olv54c, Olv56a, Olv65e, Olv67d, Olv98a, Olv98b, Van79]. Solutions [CYH22, Dun90, KYD20, OS65, AHS^{+}18, Alj21, Bal79, BD86, FBZ21, FLS14, GM60, LLS80, LF90, LOQZ14, OO94, OO95a, Olv50, Olv56b, Olv58, Olv67b, Olv75b, Olv93a, Olv93b, Olv94a, Olv97a, Olv99, Par75, PW80, S.15, SQHY08, SHY11, Sib90, Sie90, SV90a, SV90b, Waz07, Wil64, YX14]. solving [ZAAN16]. Some [Bat46, DR90, LMS73, Ojh73, Olv52b, Wim90, Won89, Cha71, FW80a, Fra81, FW85, Olv51, Sta83, Won83]. somme [Par74]. Sommerfeld [Rei90]. Sookne [HT94]. sparse [GN14]. Special [Ano11a, BCLO11, LO70, Olv72, Olv74b, Olv97b, Wim75, AGIN94, Ask90, CC85, DIW00, LLS80, LO94a, LO94b, Olv65d, Olv66, Olv75d, Olv78a, Olv90a, Tem89, Won89, Zan95]. spetsial′ [Olv78a, Olv90a]. square [CO86]. squaring [CT89]. stable [Gau90]. Staff [Ano11b]. Standards [Ano11b, Lid01]. States [Wil64]. stationary [Olv74c]. Stationery [F.63]. steepest [Kam90, MV90, Olv70b, Olv70c]. step [IKS90]. Stiefel [Sch69]. Stieltjes [SV94]. Stokes [Hin79, OO95b, OCK^{+}95, Olv90c, Par92]. structure [Shi90]. Struve [Zan95]. Student [IK76]. Sturm [Boy81, Mul73]. sub [JTB^{+}13]. sub-equation [JTB^{+}13]. subharmonic [KW90]. Subroutine [Kod08]. subspaces [DR90]. Successor [Ano11d]. Sufficient [Olv78d]. sum [Par74]. Summer [Tur89]. superintegrable [KM14]. supernonlinear [AHS^{+}18]. surface [Olv89c]. Survey [Olv63b, COT89]. Symbolic [CC85]. symmetric [CT88]. Symposia [LO70]. Symposium [Gau94, IS87, Wil64]. system [CT88, Dam83, Dam86, Olv86c, SS90, YL12]. Systems [Zah90, Gau90, KM14, LF90]. Szego [LW14, QW88].

t [Bal79]. Table [Car87, Car88, OT87]. Tables [AS64, GGM58, Has60, Olv62, Olv65a, CO55, F.63]. Tasso [AHS^{+}18, Alj21, FBZ21, JTB^{+}13, KYD20, S.15, SQHY08, SHY11, Waz07, YL12, YX14, Zay11, ZAAN16]. Taylor [Wim90]. technique [Oli68, SS90]. Technology [Lid01, Ano11b]. Temme [Olv97b]. term [BG88, Par74, Sha88]. terme [Par74]. terms [Olv80c]. th [Par74]. Their [Olv72, FLO04a, GS14, Wil64]. Theoretical [BEZHN15]. Theory [Ano11a, Bat46, Kno70, Olv65b, Rei90, Olv86e, Olv89a]. theta [Koo14]. Third [Car88, Dav78]. Thorne [UO61]. those [GS14]. three [Bha73, Sik83]. torsion [Wan90]. transform [Dai94, OR71]. Transformation [Olv49, Smi86]. transformations [Car90, SHY11]. transforms [Dai94, FW85, Olv91b, Olv14b]. transition [Hea75, Olv56a, Olv77c]. Transmission [Dam83, Dam86, Boe91, Dam83]. transmission-line [Boe91]. transparency [Dam86]. trapezoidal [AB87]. traveling [AHS^{+}18]. Treatment [SV94]. tribute [Ber14]. Trigiante [Olv89a]. tripling [KW90]. tunneling [Mey90]. Turning [Dun90, Olv63a, Olv64c, Olv86e, WH86a, WH86b, BD86, Dun14, Lak72, Olv65b, Olv75c, Olv76a, Olv77a, Olv77b, PM81]. Turning-Point [Olv63a, Olv64c]. Two [Olv61b, Bal79, Hea75, Kno70, Koo14, Olv75c, Olv76a, SFB90]. type [Bha73, FW80b, GN14, Won81].

ultraspherical [DM14]. understanding [Ask90]. Unified [SV94]. Uniform [Bal79, Boe91, BD86, Dun89, Dun90, FH90, FW88, Fre90a, Jon90, KD14, Olv58, Olv59b, Olv76b, Olv91b, Olv91c, Tem78, Tem83, WH86a, Wim80, Con90, Fre90b, LR86, Olv80c, SW88, Son80, Tem82, Won81, Won89]. Uniformly [Ski80, Ski81, Urs87]. uniqueness [CYH22, Olv99]. United [Olv91a, Wil64]. unity [Olv97a]. univalence [KT60]. Universal [GS90]. University [MW92, Olv65a, Wil64]. unknown [Olv61c]. Unrestricted [CO86, Olv80b, Olv85, CO80]. Unsolved [Olv75d]. upper [EL98]. Ursell [MW92, Mac75, SS90]. USA [SKL93]. Use [SL72, CO55]. Using [OT87, CT89, FLO04a, HC91, Tem89].

V [Jor80a, Olv89a, F.63, Olv78a, Olv84]. VA [SKL93]. Valid [Ski80, Ski81]. value [AO90, Lak72]. values [FW80a, Fra81, Olv54c, Olv58, Olv60]. Vancouver [Gau94]. Vandermonde [Gau90]. vanishingly [Bal79]. variable [Kam90, Olv14b, Sha88]. variables [Par90]. various [FBZ21, Ism77]. Vectors [Pry84]. Venerable [Ano11d]. Via [SV94, Fre90b, Smi86]. viewed [AO90]. Villa [IS87]. Vol [Sch69]. Vols [Olv72]. Volterra [SV94]. Volume [Olv65a]. Vvedenie [Olv78a].

W

[AGIN94, Ano13b, BB13, Cro11, F.63, Tho11, Wim75, Won90, Won00, Won13, Won14b]. Washington [LO70]. Wasow [Olv86e]. Wave [LO70, MW92, Olv65a, Dav90, FBZ21, LLS80, MV90, Olv65c, SQHY08]. wave-penetration [Olv65c]. wave-propagation [MV90]. Waves [BEZHN15, AHS^{+}18]. Weber [Olv59b]. Whittaker [Dun89, Olv65e, Olv80c]. whose [Mul73, Par74]. Wiley [Olv91a]. Wimp [Olv86a]. Wisconsin [Wil64]. Wissenschaften [Sch69]. without [Hay09, Olv83c]. WKB [Olv61a]. Wolfgang [Olv86e]. workshop [DIW00]. Wrigge [Fra81].

xxx [FLO04b]. xxxv [Olv65a].

York [Olv65a, Olv89a, Olv91a]. Yudell [Olv72].

Zero [DV14]. Zeros [DM14, Olv60, Bot77, EL98, FO99, FLO04a, Gat90, GS14, HB75, McC77, Mul79, Olv50, Olv51, Olv52a, Par74, SP73, SV90a, SV90b, VRZG96, Par74, TL76]. zeta [Has60]. zillion [Hay09]. Zur [Ber77].

References

Allasia:1987:NCI

[AB87] Giampietro Allasia and Renata Besenghi. Numerical calculation of incomplete gamma functions by the trapezoidal rule. Numerische Mathematik, 50(4): 419–428, July 1987. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Andrews:1994:D

[AGIN94] George Andrews, George Gasper, Mourad Ismail, and Paul Nevai. Dedication: [special issue to Richard A. Askey and Frank W. J. Olver]. SIAM Journal on Mathematical Analysis, 25 (2):vii–ix, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL https://epubs.siam.org/
toc/sjmaah/25/2.

Ali:2018:ESC

[AHS^{+}18] Muhammad Nasir Ali, Syed Muhammad Husnine, Asit Saha, Samir Kumar Bhowmik, Sharanjeet Dhawan, and Turgut Ak. Exact solutions, conservation laws, bifurcation of nonlinear and supernonlinear traveling waves for Sharma– Tasso–Olver equation. Nonlinear Dynamics, 94(3):1791–1801, July 2018. CODEN NODYES. ISSN 0924-090X (print), 1573-269X (electronic).

Aljoudi:2021:ESF

[Alj21] Shorog Aljoudi. Exact solutions of the fractional Sharma–Tasso– Olver equation and the fractional Bogoyavlenskii’s breaking soliton equations. Applied Mathematics and Computation, 405(??): Article 126237, September 15, 2021. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S0096300321003271.

Anonymous:2008:NRP

[Ano08]  Anonymous.  NIST releases

preview of much-anticipated online mathematics reference. NIST Web site, June 24,

2008. URL https://www.nist.
gov/news-events/news/2008/
06/nist-releases-preview-much-

anticipated-online-mathematics-
reference.

Anonymous:2010:DLM

[Ano10a] Anonymous. Digital Library of Mathematical Functions. NIST Web site, 2010. URL https:

//www.nist.gov/mathematics-
statistics/digital-library-
mathematical-functions.

Anonymous:2010:PGHa

[Ano10b] Anonymous. Portrait gallery 2010 honorees. NIST Web site, October 4, 2010. URL https://

www.nist.gov/director/portrait-
gallery-2010-honorees. Frank

W. J. Olver (ITL): Years at NBS/NIST:1961–1986. For prolific contributions in applied mathematics, including those to the NBS Handbook of Mathematical Functions (1964) and its successor, the NIST Digital Library of Mathematical Functions (2010).

Anonymous:2011:ICS

[Ano11a] Anonymous. International conference on special functions in the 21st century: Theory and applications. NIST Web site, April 6–8, 2011.

URL https://www.nist.gov/
news-events/events/2011/04/
international-conference-special-
functions-21st-century-theory-

and. This conference is dedicated to Professor Olver in light of his seminal contributions to the advancement of special functions, especially in the area of asymptotic analysis and as Mathematics Editor of the DLMF.

Anonymous:2011:NIS

[Ano11b] Anonymous. National Institute of Standards and Technology recognizes staff. NIST Web site, December 7, 2011. URL https:

                                        //www.nist.gov/news-events/
                                        news/2011/12/national-institute-
                                        standards-and-technology-recognizes-
                                        staff.

Anonymous:2011:NDL

[Ano11c] Anonymous. NIST Digital Library of Mathematical Functions receives IT award. NIST Web site, October 25,

                                        2011. URL https://www.nist.
                                        gov/news-events/news/2011/
                                        10/nist-digital-library-mathematical-
                                        functions-receives-it-award.

Anonymous:2011:NRS

[Ano11d] Anonymous. NIST releases successor to venerable Handbook of Math Functions. NIST Web site, May 11, 2011. URL https:

                                        //www.nist.gov/news-events/
                                        news/2010/05/nist-releases-
                                        successor-venerable-handbook-
                                        math-functions.

Anonymous:2010:PGHb

[Ano11e] Anonymous. Portrait gallery 2011 honorees. NIST Web site, October 26, 2011. URL https://

                                        www.nist.gov/director/portrait-
                                        gallery-honorees-2011. Frank

W. J. Olver (ITL): Years at NBS/NIST: 1961–1986: For prolific contributions in applied mathematics, including those to the NBS Handbook of Mathematical Functions (1964) and its successor, the NIST Digital Library of Mathematical Functions (2010).

Anonymous:2013:FOO

[Ano13a] Anonymous. Frank Olver obituary. Web site, April 26, 2013.

URL https://www.legacy.com/
us/obituaries/washingtonpost/
name/frank-olver-obituary?
id=6008137.

Anonymous:2013:FWJ

[Ano13b] Anonymous. Frank W. J.

Olver, mathematician. Washington Post, ??(??):??, April 30, 2013. ISSN 0190-8286. URL

https://www.washingtonpost.
com/local/obituaries/frank-
wj-olver-mathematician/2013/
04/30/51080646-b1b0-11e2-baf7-
5bc2a9dc6f44_story.html.

Allen:1990:SPB

[AO90] J. D. Allen and Robert E.

O’Malley, Jr. Singularly perturbed boundary value problems viewed in the Liénard plane. In Wong [Won90], pages 357– 378. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Abramowitz:1964:HMF

[AS64] Milton Abramowitz and Irene A.

Stegun, editors. Handbook of Mathematical Functions with Formulas, Graphs, and Mathematical Tables, volume 55 of Applied mathematics series. U. S. Department of Commerce, Washington, DC, USA, 1964. xiv + 1046 pp. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://www.

convertit.com/Go/ConvertIt/
Reference/AMS55.ASP in bitmap

image format.

Askey:1990:GAU

[Ask90] Richard A. Askey. Graphs as an aid to understanding special functions. In Wong [Won90], pages 3–33. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Baldwin:1979:UAS

[Bal79] P. Baldwin. Uniform approximations to solutions of a linear second-order differential equation outside a vanishingly small region containing two asymptotically coincident t. Quarterly Journal of Mechanics and Applied Mathematics, 32(2):187–204, 1979. CO-DEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Bateman:1946:SIE

[Bat46] Harry Bateman. Some integral equations of potential theory. Journal of Applied Physics, 17 (2):91–102, February 1946. CO-DEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic),

1520-8850. URL https://pubs.
aip.org/aip/jap/article/17/
2/91/158466/Some-Integral-
Equations-of-Potential-Theory.

See corrections [Olv46].

Bailey:2013:FWJ

[BB13] David H. Bailey and Jonathan M.

Borwein. Frank W. J. Olver (1924–2013). Math Drudge, May 12, 2013. URL https://

experimentalmath.info/blog/
2013/05/frank-w-j-olver-1924-
2013/.

Boisvert:2011:SFH

[BCLO11] Ronald F. Boisvert, Charles W.

Clark, Daniel W. Lozier, and Frank W. Olver. A special functions handbook for the digital age. Notices of the American Mathematical Society, 58(7): 905–911, August 1, 2011. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

URL https://www.ams.org/
journals/notices/201107/201107-
full-issue.pdf;   https://
www.nist.gov/publications/
special-functions-handbook-
digital-age.

Boyd:1986:UAS

[BD86] W. G. C. Boyd and T. M. Dunster. Uniform asymptotic solutions of a class of second-order linear differential equations having a turning point and a regular singularity, with an application to Legendre functions. SIAM Journal on Mathematical Analysis, 17(2):422–450, March 1986. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Berndt:1984:CRS

[BE84] Bruce C. Berndt and Ronald J.

Evans. Chapter 13 of Ramanujan’s second notebook: integrals and asymptotic expansions. Exposition. Math., 2(4):289–347, 1984. ISSN 0723-0869 (print), 1878-0792 (electronic).

Berndt:1992:AES

[BE92] Bruce C. Berndt and Ronald J. Evans. Asymptotic expansion of a series of Ramanujan. Proceedings of the Edinburgh Mathematical Society (2), 35(2):189– 199, 1992. CODEN PEMSA3. ISSN 0013-0915 (print), 1464-3839 (electronic).

Berg:1977:ARA

[Ber77]  L. Berg.  Zur Abschätzung

des Restgliedes in der asymptotischen Entwicklung des Exponentialintegrals. Computing, 18(4):361–363, 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Berg:1990:AIS

[Ber90] L. Berg. Asymptotics of integrals, series, and operators. In Wong [Won90], pages 35– 52. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Berry:2014:TFO

[Ber14] Michael Berry. A tribute to Frank Olver (1924–2013). Analysis and Applications (Singapore), 12(4): ix–x, June 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part I).

Belafhal:2015:TIG

[BEZHN15] Abdelmajid Belafhal, Lahcen Ez-Zariy, Salima Hennani, and Hamid Nebdi. Theoretical introduction and generation method of a novel nondiffracting waves: Olver beams. Optics and Photonics Journal, 05(07):234–246, 2015. CODEN ???? ISSN 2160-8881 (print), 2160-889X (electronic).

Baratella:1988:BET

[BG88] Paola Baratella and Luigi Gatteschi. The bounds for the error term of an asymptotic approximation of Jacobi polynomials. Lecture Notes in Mathematics, 1329:203–221, 1988. CODEN LNMAA2. ISBN 3-540-19489-4 (print), 3-540-39295-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0083360/.

Berry:1990:H

[BH90] M. V. Berry and C. J. Howls.

Hyperasymptotics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 430(1880):653– 668, 1990. CODEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Bhatnagar:1973:LTI

[Bha73] P. L. Bhatnagar. Lommel type integrals involving products of three Bessel functions. Comment. Math. Univ. St. Paul., 22(1):1– 11, 1973. ISSN 0010-258X.

Bauer:1967:DA

[BHO^{+}67] Friedrich L. Bauer, Alston S. Householder, Frank W. J. Olver, Heinz Rutishauser, Klaus Samelson, and Eduard Stiefel. Description of ALGOL 60, volume 1a of Handbook for Automatic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-540-03826-4, 3-642-86934-3 (ebook), 3-642-86936-X (print), 3-662-38103-6. xii + 326 pp. LCCN QA76.5 .R87. URL http://www.

springerlink.com/content/978-
3-642-86934-1.

Boisvert:2001:HMF

[BL01] Ronald F. Boisvert and Daniel W. Lozier. Handbook of Mathematical Functions. In Lide [Lid01], pages 135–139. URL https:/

/nvlpubs.nist.gov/nistpubs/
sp958-lide/135-139.pdf;   https:
//nvlpubs.nist.gov/nistpubs/
sp958-lide/html/135-139.html.

NIST Special Publication.

Boersma:1976:PLI

[Boe76]  J. Boersma.  Problem 75-

20, “Limit of an integral”, by M. L. Glasser. SIAM Review, 18(4):770–772, 1976. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Boersma:1991:UAB

[Boe91] J. Boersma. Uniform asymptotics of a Bessel-function series occurring in a transmission-line problem. Journal of Computational and Applied Mathematics, 37(1–3):143–159, 1991. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Bottema:1977:ZBP

[Bot77] O. Bottema. On the zeros of the Bessel polynomials. Indagationes Mathematicæ, 39(5):380– 382, 1977. CODEN IMTHBJ. ISSN 0019-3577, 0023-3358. Nederl. Akad. Wetensch. Proc. Ser. A 80.

Boyd:1981:SLE

[Boy81] John P. Boyd. Sturm–Liouville eigenproblems with an interior pole. Journal of Mathematical Physics, 22(8):1575–1590, August 1981. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.
org/resource/1/jmapaq/v22/
i8/p1575_s1.

Boyd:1990:AEC

[Boy90] W. G. C. Boyd. Asymptotic expansions for the coefficient functions associated with linear second-order differential equations: the simple pole case. In Wong [Won90], pages 53– 73. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Baricz:2014:RCN

[BS14] Árpád Baricz and Róbert Szász.

The radius of convexity of normalized Bessel functions of the first kind. Analysis and Applications (Singapore), 12(5):485–509, August 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part II).

Besieris:2019:NSA

[BSD19] Ioannis M. Besieris, Amr M. Shaarawi, and Bradley Davis. A note on self-accelerating Olver and Olver–Gauss beams. Optics and Photonics Journal, 09(01):1– 7, 2019. ISSN 2160-889X.

Buschman:1976:LFI

[Bus76] R. G. Buschman. Legendre functions, inequalities and partial derivatives. Ranchi Univ. Math. J., 7:28–32, 1976. ISSN 0079-9602.

Carlson:1987:TEI

[Car87] B. C. Carlson. A table of elliptic integrals of the second kind. Mathematics of Computation, 49 (180):595–606, S13–S17, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1988:TEI

[Car88] B. C. Carlson. A table of elliptic integrals of the third kind. Mathematics of Computation, 51(183): 267–280, S1–S5, July 1988. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1990:LTI

[Car90] B. C. Carlson. Landen transformations of integrals. In Wong [Won90], pages 75–94. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Cash:1978:EOM

[Cas78]  J. R. Cash.  An extension

of Olver’s method for the numerical solution of linear recurrence relations. Mathematics of Computation, 32(142):497– 510, April 1978. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cash:1980:NOA

[Cas80]  J. R. Cash.  A note on

Olver’s algorithm for the solution of second-order linear difference equations. Mathematics of Computation, 35(151): 767–772, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL https://www.ams.org/
journals/mcom/1980-35-151/
S0025-5718-1980-0572854-5.

Calmet:1985:SMR

[CC85] J. Calmet and I. Cohen. Symbolic manipulation of recurrence relations. An approach to the manipulation of special functions. In The second RIKEN international symposium on symbolic and algebraic computation by computer (Wako-shi, 1984), volume 2 of World Sci. Ser. Comput. Sci., pages 55–65. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1985. ISBN 9971-5-0021-3.

Ciasullo:1990:ACC

[CC90] L. M. Ciasullo and James A. Cochran. Accelerating the convergence of Chebyshev series. In Wong [Won90], pages 95– 136. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Carinhas:1990:CAF

[CF90] P. A. Carinhas and S. A. Fulling. Computational asymptotics of fourth-order operators. In Wong [Won90], pages 601–617. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Clenshaw:1961:MCM

[CGM^{+}61] C. W. Clenshaw, E. T. Goodwin, D. W. Martin, G. F. Miller, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of National Physical Laboratory. Notes on Applied Science. Her Majesty’s Stationery Office, London, UK, second edition, 1961. vi + 170 pp. LCCN QA297 .T4 1961.

Cox:1990:RNC

[CH90] M. G. Cox and S. Hammarling, editors. Reliable numerical computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Chatterjea:1971:ODS

[Cha71] S. K. Chatterjea. Operational derivation of some generating functions of the Bessel polynomials. Math. Balkanica, 1:293–297, 1971. ISSN 0350-2007.

Chen:1991:AJP

[CI91] Li-Chen Chen and Mourad E. H.

Ismail. On asymptotics of Jacobi polynomials. SIAM Journal on Mathematical Analysis, 22(5): 1442–1449, September 1991. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Clenshaw:1986:GEL

[CLOT86] C. W. Clenshaw, Daniel W.

Lozier, F. W. J. Olver, and Peter R. Turner. Generalized exponential and logarithmic functions. Computers and Mathematics with Applications. Part B, 12 (5–6):1091–1101, 1986. CODEN CMAPDK. ISSN 0886-9561.

Clenshaw:1951:NNA

[CO51] C. W. Clenshaw and F. W. J.

Olver. Notes on numerical analysis—3: Solution of differential equations by recurrence relations (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 5(33):34–39, January 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clenshaw:1955:UEP

[CO55] C. W. Clenshaw and Frank W. J. Olver. The use of economized polynomials in mathematical tables. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 51(4):614–628, October 1955. CODEN PCPSA4. ISSN 0008-1981.

Clenshaw:1980:UAE

[CO80] C. W. Clenshaw and Frank W. J. Olver. An unrestricted algorithm for the exponential function. SIAM Journal on Numerical Analysis, 17(2):310– 331, 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Clenshaw:1984:BFP

[CO84] C. W. Clenshaw and Frank W. J. Olver. Beyond floating point. Journal of the ACM, 31(2):319– 328, 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Clenshaw:1986:UAR

[CO86] C. W. Clenshaw and Frank W. J. Olver. Unrestricted algorithms for reciprocals and square roots. BIT (Nordisk tidskrift for informationsbehandling), 26 (4):476–492, 1986. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Clenshaw:1987:LIA

[CO87] C. W. Clenshaw and Frank W. J. Olver. Level-index arithmetic operations. SIAM Journal on Numerical Analysis, 24(2):470– 485, 1987. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coleman:1980:DRO

[Col80] J. P. Coleman. Divergence rules OK? Mathematical Gazette, 64 (429):180–184, 1980. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Connor:1990:PMU

[Con90] J. N. L. Connor. Practical methods for the uniform asymptotic evaluation of oscillating integrals with several coalescing saddle points. In Wong [Won90], pages 137–173. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Clenshaw:1989:LIA

[COT89] C. W. Clenshaw, F. W. J. Olver, and Peter R. Turner. Levelindex arithmetic: an introductory survey. In Numerical analysis and parallel processing (Lancaster Numerical Analysis Summer School, 1987), volume 1397 of Lecture Notes in Math., pages 95–168. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51645-X.

Crowder:2011:SBRa

[Cro11] Martin Crowder. Short book review: NIST Handbook of Mathematical Functions edited by Frank W. J. Olver, Daniel W. Lozier, Ronald F. Boisvert, Charles W. Clark. International Statistical Review = Revue Internationale de Statistique, 79 (1):131–132, April 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

                                        URL http://www.jstor.org/
                                        stable/41306184.

Clenshaw:1988:SLI

[CT88] C. W. Clenshaw and Peter R. Turner. The symmetric levelindex system. IMA Journal of Numerical Analysis, 8(4):517– 526, October 1988. CODEN IJ-NADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Clenshaw:1989:RSU

[CT89] C. W. Clenshaw and Peter R. Turner. Root squaring using level-index arithmetic. Computing, 43(2):171–185, June 1989. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Cohl:2014:EFS

[CV14] Howard S. Cohl and Hans Volkmer. Expansions for a fundamental solution of Laplace’s equation on R^{3} in 5-cyclidic harmonics. Analysis and Applications (Singapore), 12(6):613–633, October 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part III).

Chen:2022:PSN

[CYH22] Aiyong Chen, Hao Yu, and Xiaokai He. 2-peakon solutions and non-uniqueness of the Fokas– Olver–Rosenau–Qiao equation with higher-order nonlinearity. Bulletin of the Malaysian Mathematical Sciences Society, 45 (6):3401–3422, September 2022. ISSN 0126-6705 (print), 2180-4206 (electronic).

Dai:1994:GAE

[Dai94] Hui-Hui Dai. On the general asymptotic expansions of H_{v}^{(1)}- transform and related Bessel transforms. Analysis, 14(1):19– 41, 1994. ISSN 0174-4747.

Dammert:1983:TTS

[Dam83] Örjan Dammert. Transmission through a system of potential barriers. I. Transmission coefficient. Journal of Mathematical Physics, 24(8):2163–2175, August 1983. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.
org/resource/1/jmapaq/v24/
i8/p2163_s1.

Dammert:1986:TTS

[Dam86] Örjan Dammert. Transmission through a system of potential barriers. II. Necessary condition for complete transparency. A maximum transmission problem. Journal of Mathematical Physics, 27(2):461–470, February 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.
org/resource/1/jmapaq/v27/
i2/p461_s1.

Davis:1978:NEL

[Dav78] A. M. J. Davis. Non-existence of large eigenvalues of a third order differential equation. Mathematika, 25(1):151–159, 1978. CO-DEN MTKAAB. ISSN 0025-5793.

Davis:1990:CSW

[Dav90] Anthony M. J. Davis. Continental shelf wave scattering: partial removal of the rigid lid. In Wong [Won90], pages 555–563. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Dunkl:2000:PIW

[DIW00] Charles F. Dunkl, Mourad Ismail, and Roderick Wong, editors. Proceedings of the international workshop, special functions: Hong Kong, 21–25 June 1999. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2000. ISBN 981-02-4393-6. LCCN QA351.P76 2000; QC20.7.F87 P76 2000.

Dunster:1991:CFS

[DL91] T. M. Dunster and D. A. Lutz. Convergent factorial series expansions for Bessel functions. SIAM Journal on Mathematical Analysis, 22(4):1156–1172, July 1991. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Driver:2014:ZPU

[DM14] Kathy Driver and Martin E.

Muldoon. Zeros of pseudoultraspherical polynomials. Analysis and Applications (Singapore), 12(5):563–581, August 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part II).

Dieci:1990:SAI

[DR90] Luca Dieci and Robert D. Russell. Some aspects of invariant subspaces computation. In Wong [Won90], pages 565–585. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Dunster:1989:UAE

[Dun89] T. M. Dunster. Uniform asymptotic expansions for Whittaker’s confluent hypergeometric functions. SIAM Journal on Mathematical Analysis, 20(3):744–760, May 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:1990:UAS

[Dun90] T. M. Dunster. Uniform asymptotic solutions of second-order linear differential equations having a double pole with complex exponent and a coalescing turning point. SIAM Journal on Mathematical Analysis, 21(6):

1594–1618, November 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Dunster:1991:CFO

[Dun91] T. M. Dunster. Conical functions with one or both parameters large. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and Physical Sciences, 119(3–4):311–327, 1991. CODEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic).

Dunster:2014:OEB

[Dun14] T. M. Dunster. Olver’s error bound methods applied to linear ordinary differential equations having a simple turning point. Analysis and Applications (Singapore), 12(4):385–402, June 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part I).

Duval:1983:EAI

[Duv83] Anne Duval. Étude asymptotique d’une intégrale analogue à la fonction ‘Γ modifiée’. (French) []. Lecture Notes in Mathematics, 1015:50–63, 1983. CODEN LNMAA2. ISBN 3-540-12684-8 (print), 3-540-38674-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0071349/;   https:
//www.math.utah.edu/pub/bibnet/
authors/o/olver-frank-w-j.
bib.

Dominici:2014:ZDP

[DV14] Diego Dominici and Walter Van Assche. Zero distribution of polynomials satisfying a differentialdifference equation. Analysis and Applications (Singapore), 12(6):635–666, October 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part III).

Elbert:1998:UBZ

[EL98] Árpád Elbert and Andrea Laforgia. An upper bound for the zeros of the cylinder function C_{ν}(x). Mathematical Inequalities & Applications, 1(1):105–111, 1998. ISSN 1331-4343 (print), 1848-9966 (electronic).

F:1963:BRF

[F.63] A. F. Book review: F. W. J.

Olver, Tables for Bessel Functions of Moderate or Large Orders, v. 6, Her Majesty’s Stationery Office, London, 1962, iii + 51 pp. Mathematics of Computation, 17(84):464–466, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fabijonas:2004:AAF

[Fab04] B. R. Fabijonas. Algorithm 838: Airy functions. ACM Transactions on Mathematical Software, 30(4):491–501, December 2004. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/10.
1145/1039813.1039819.

Feng:2021:EBV

[FBZ21] Y. Y. Feng, S. D. Bilige, and R. F. Zhang. Evolutionary behavior of various wave solutions of the (2 + 1)-dimensional Sharma– Tasso–Olver equation. Indian Journal of Physics, 96(7):2107– 2114, July 2021. CODEN IJP-NCV. ISSN 0973-1458 (print), 0974-9845 (electronic).

Fox:1957:MCM

[FGM^{+}57] L. Fox, E. T. Goodwin, J. G. L. Michel, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of Notes on applied science. Her Majesty’s Stationery Office, London, UK, 1957. vi + 128 pp. LCCN QA297 .T4.

Fox:1970:MCM

[FGM^{+}70] L. Fox, E. T. Goodwin, J. G. L. Michel, F. W. J. Olver, and J. H. Wilkinson. Modern Computing Methods, volume 16 of NPL Notes on Applied Science. Her Majesty’s Stationery Office, London, UK, second edition, 1970. ISBN 0-11-480021-9. vii + 169 pp. LCCN QA297 .T4 1970.

Fitouhi:1990:UEE

[FH90] A. Fitouhi and M. M. Hamza. A uniform expansion for the eigenfunction of a singular secondorder differential operator. SIAM Journal on Mathematical Analysis, 21(6):1619–1632, November 1990. CODEN SJMAAH.

ISSN 0036-1410 (print), 1095-7154 (electronic).

Fabijonas:2004:CCA

[FLO04a] B. R. Fabijonas, Daniel W.

Lozier, and F. W. J. Olver. Computation of complex Airy functions and their zeros using asymptotics and the differential equation. ACM Transactions on Mathematical Software, 30(4): 471–490, 2004. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fabijonas:2004:AXA

[FLO04b] Bruce R. Fabijonas, Daniel W.

Lozier, and Frank W. Olver. Algorithm xxx: Airy functions. NIST Web site, December 15,

2004. URL https://www.nist.
gov/publications/algorithm-
xxx-airy-functions.

Ferreira:2014:CAE

[FLS14] Chelo Ferreira, José L. López, and Ester Pérez Sinusía. Convergent and asymptotic expansions of solutions of second-order differential equations with a large parameter. Analysis and Applications (Singapore), 12(5):523–536, August 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part II).

Fabijonas:1999:RAE

[FO99] Bruce R. Fabijonas and F. W. J.

Olver. On the reversion of an asymptotic expansion and the zeros of the Airy functions. SIAM Review, 41(4):762– 773, 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Fransen:1981:ACH

[Fra81] Arne Fransén. Addendum and corrigendum to: “High-precision values of the gamma function and of some related coefficients” [Math. Comp. 34 (1980), no. 150, 553–566, MR 81f:65004] by Fransén and S. Wrigge. Mathematics of Computation, 37(155): 233–235, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [FW80a].

Frenzen:1990:EBU

[Fre90a] C. L. Frenzen. Error bounds for a uniform asymptotic expansion of the Legendre function Q_{n}^{−m}(cosh z). SIAM Journal on Mathematical Analysis, 21 (2):523–535, March 1990. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Frenzen:1990:EBC

[Fre90b] C. L. Frenzen. Error bounds via complete monotonicity for a uniform asymptotic expansion of the Legendre function P_{n}^{−m}(cosh z). In Wong [Won90], pages 587– 599. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Fransen:1980:HPV

[FW80a] Arne Fransén and Staffan Wrigge. High-precision values of the gamma function and of some related coefficients. Mathematics of Computation, 34(150):553–566, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See addendum and corrigendum [Fra81].

Fu:1980:AEB

[FW80b] J. C. Fu and R. Wong. An asymptotic expansion of a betatype integral and its application to probabilities of large deviations. Proceedings of the American Mathematical Society, 79(3): 410–414, 1980. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Frenzen:1985:NAE

[FW85] C. L. Frenzen and R. Wong. A note on asymptotic evaluation of some Hankel transforms. Mathematics of Computation, 45(172): 537–548, October 1985. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frenzen:1988:UAE

[FW88] C. L. Frenzen and R. Wong.

Uniform asymptotic expansions of Laguerre polynomials. SIAM Journal on Mathematical Analysis, 19(5):1232–1248, September 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gatteschi:1990:NIZ

[Gat90] Luigi Gatteschi. New inequalities for the zeros of confluent hypergeometric functions. In Wong [Won90], pages 175–192.

ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Gautschi:1990:HSV

[Gau90] Walter Gautschi. How (un)stable are Vandermonde systems? In Wong [Won90], pages 193–210. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943– 1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9– 13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gupta:1958:TP

[GGM58] Hansraj Gupta, C. E. Gwyther, and J. C. P. Miller. Tables of Partitions, volume 4 of Royal Society Mathematical Tables. Cambridge University Press, Cambridge, UK, 1958. xxxix + 132 pp.

Gupta:1960:RPQ

[GM60] Hansraj Gupta and Jeffery Charles Percy Miller, editors. Representations of primes by quadratic forms, displaying solutions of the Diophantine equation kp = a^{2} + Db^{2}, volume 5 of Royal Society Mathematical Tables. Cambridge University Press, Cambridge, UK, 1960. 135 pp.

Gawronski:2014:CSB

[GN14] Wolfgang Gawronski and Thorsten Neuschel. On a conjecture on sparse binomial-type polynomials by Brown, Dilcher and Manna. Analysis and Applications (Singapore), 12(5):511–522, August 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part II).

Glasser:1977:ABI

[GO77] M. L. Glasser and F. W. J. Olver.

Asymptotic behaviour of integrals of Bessel functions of high order. Utilitas Mathematica, 12: 225–239, 1977. CODEN UT-MADA. ISSN 0315-3681. URL

https://combinatorialpress.
com/um/arch/vol12/.

Goh:1990:UAD

[GS90] William M. Y. Goh and Eric Schmutz. Universal asymptotic distribution functions mod 1. In Wong [Won90], pages 619–626. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Gil:2014:CZA

[GS14] Amparo Gil and Javier Segura. On the complex zeros of Airy and Bessel functions and those of their derivatives. Analysis and Applications (Singapore), 12(5): 537–561, August 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part II).

Hanson:1990:SPE

[Han90]  Floyd B. Hanson.  Singular

point and exponential asymptotics. In Wong [Won90], pages 211–240. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Haselgrove:1960:TRZ

[Has60] C. B. Haselgrove, editor. Tables of the Riemann zeta function, volume 6 of Royal Society Mathematical Tables. Cambridge University Press, Cambridge, UK, 1960. xxii + 80 pp. LCCN QA351 .H3.

Hayes:2009:HA

[Hay09] Brian Hayes. The higher arithmetic: How to count to a zillion without falling off the end of the number line. American Scientist, 97(5):364, 2009. CODEN AM-SCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Hayes:2017:FOM

[Hay17] Brian Hayes. Foolproof, and Other Mathematical Meditations.

MIT Press, Cambridge, MA, USA, 2017. ISBN 0-262-03686-X. x + 234 pp. LCCN QA93 .H3586 2017.

Headley:1975:DZG

[HB75] V. B. Headley and V. K. Barwell.

On the distribution of the zeros of generalized Airy functions. Mathematics of Computation, 29 (131):863–877, July 1975. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hsu:1991:KAE

[HC91] Leetsch C. Hsu and Wen Chang Chu. A kind of asymptotic expansion using partitions. Tohoku Math. J. (2), 43(2):235–242, 1991. ISSN 0040-8735.

Heading:1972:FEA

[Hea72] John Heading. Further exact and approximate considerations of the barrier problem. Journal of the Institute of Mathematics and its Applications, 10(??):312–324, 1972. CODEN JMTAA8. ISSN 0020-2932.

Heading:1975:BBT

[Hea75] John Heading. Barriers bounded by two transition points of arbitrary odd order. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and Physical Sciences, 73:51– 64, 1975. CODEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic).

Hinton:1979:SMC

[Hin79] F. L. Hinton. Stokes multipliers for a class of ordinary differential equations. Journal of Mathematical Physics, 20(10): 2036–2046, October 1979. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/
v20/i10/p2036_s1.

Hasegawa:1994:EOS

[HT94] Takemitsu Hasegawa and Tatsuo Torii. An extension of the Olver– Sookne method for the solution of second-order linear difference equations. In Gautschi [Gau94], pages 297–300. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. URL http:/

/math.nist.gov/nesf/. See also

SIAM Review, September 1995, 37(3), p. 483.

Igarashi:1985:PPA

[Iga85] Masao Igarashi. Practical problems arising for finding roots of nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 1(5):433–455, September 1985. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ismail:1976:BPS

[IK76] Mourad E. H. Ismail and Douglas H. Kelker. The Bessel polynomials and the Student t distribution. SIAM Journal on Mathematical Analysis, 7(1):82– 91, 1976. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ikebe:1990:RAS

[IKS90] Yasuhiko Ikebe, Marek Kowalski, and Frank Stenger. Rational approximation of the step, filter, and impulse functions. In Wong [Won90], pages 441–454. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Irwin:1987:PSC

[IS87] Mary Jane Irwin and Renato Stefanelli, editors. Proceedings: 8th Symposium on Computer Arithmetic, May 19–21, 1987, Villa Olmo, Como, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 1987. ISBN 0-8186-0774-2 (paperback), 0-8186-4774-4 (microfiche), 0-8186-8774-6 (case). LCCN QA 76.9 C62 S95 1987.

Ismail:1977:IRC

[Ism77] Mourad E. H. Ismail. Integral representations and complete monotonicity of various quotients of Bessel functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(6):1198–1207, 1977. CODEN CJMAAB. ISSN 0008-414X.

Jones:1990:UAR

[Jon90] D. S. Jones. Uniform asymptotic remainders. In Wong [Won90], pages 241–264. ISBN 0-8247-8347-6. LCCN QA299.6 .A88

1990. URL http://www.loc.

                                        gov/catdir/enhancements/fy0647/
                                        90002810-d.html.

Jones:1993:ASR

[Jon93] D. S. Jones. Asymptotic series and remainders. In Ordinary and partial differential equations, Vol. IV (Dundee, 1992), volume 289 of Pitman Res. Notes Math. Ser., pages 112–125. Longman Scientific and Technical, Harlow, Essex, 1993. ISBN 0-582-09137-3.

Jordan:1980:BRC

[Jor80a] W. B. Jordan. Book review: A Conjectured Property of Legendre Functions (A. K. Raina and V. Singh). SIAM Review, 22(3):369–372, July 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jordan:1980:CPL

[Jor80b]  W. B. Jordan.  A conjec-

tured property of Legendre functions. SIAM Review, 22(3):369– 372, 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Jafari:2013:FSE

[JTB^{+}13] Hossein Jafari, Haleh Tajadodi, Dumitru Baleanu, Abdulrahim Al-Zahrani, Yahia Alhamed, and Adnan Zahid. Fractional sub-equation method for the fractional generalized reaction Duffing model and nonlinear fractional Sharma–Tasso–Olver equation. Open Physics, 11(10), January 2013. CODEN OPPHBI. ISSN 2391-5471.

Kaminski:1990:VSP

[Kam90] D. Kaminski. On the nvariable saddle point and steepest descent methods. In Wong [Won90], pages 627–637. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Khwaja:2014:UAE

[KD14] S. Farid Khwaja and A. B. Olde Daalhuis. Uniform asymptotic expansions for hypergeometric functions with large parameters IV. Analysis and Applications (Singapore), 12(6):667–710, October 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part III).

Kalnins:2014:QAC

[KM14] Ernest G. Kalnins and W. Miller.

Quadratic algebra contractions and second-order superintegrable systems. Analysis and Applications (Singapore), 12(5):583–612, August 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part II).

Knol:1970:GTR

[Kno70] Wolter Knol. Generalizations of two relations in Bessel function theory. Rijksuniversiteit te Groningen, Groningen, The Netherlands, 1970. iii + 118 pp. Doctoral dissertation, University of Groningen.

Kodama:2008:ASP

[Kod08] Masao Kodama. Algorithm 877: A subroutine package for cylindrical functions of complex order and nonnegative argument. ACM Transactions on Mathematical Software, 34(4):22:1– 22:21, July 2008. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kohno:1979:EAF

[Koh79] Mitsuhiko Kohno. An extended Airy function of the first kind. Hiroshima Math. J., 9(2):473– 489, 1979. ISSN 0018-2079.

                                        URL http://projecteuclid.
                                        org/euclid.hmj/1206134896.

Koornwinder:2014:ETF

[Koo14]  Tom H. Koornwinder.  On

the equivalence of two fundamental theta identities. Analysis and Applications (Singapore), 12(6):711–725, October 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part III).

Kreyszig:1960:RUB

[KT60] Erwin Kreyszig and John Todd. The radius of univalence of Bessel functions. I. Illinois Journal of Mathematics, 4(2):143–149, March 1960. CODEN IJMTAW. ISSN 0019-2082 (print), 1945-6581 (electronic). URL http:

                                        //projecteuclid.org/euclid.
                                        ijm/1255455740.

Kazarinoff:1990:PTS

[KW90] Nicholas D. Kazarinoff and Joseph S. Wilkowski. Period tripling and subharmonic oscillations in Marangoni flows in a cylindrical liquid bridge. In Wong [Won90], pages 265–283. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Kaya:2020:CEN

[KYD20] Doğan Kaya, Asıf Yokuş, and Uğur Demiroğlu. Comparison of exact and numerical solutions for the Sharma–Tasso– Olver equation. In Nonlinear Systems and Complexity, pages 53– 65. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020.

Lakin:1972:BVP

[Lak72] W. D. Lakin. Boundary value problems with a turning point. Studies in Applied Mathematics, 51:261–275, 1972. CO-DEN SAPMB6. ISSN 0022-2526 (print), 1467-9590 (electronic).

Leung:1990:PSD

[LF90] Anthony W. Leung and Guangwei Fan. Positive solutions for degenerate and nondegenerate elliptic systems: existence and numerical approximations. In Wong [Won90], pages 285–298. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/

                                        enhancements/fy0647/90002810-
                                        d.html.

Lide:2001:CEM

[Lid01] D. R. Lide, editor. A Century of Excellence in Measurements, Standards, and Technology: A Chronicle of Selected NBS/NIST Publications, 1901– 2000, volume 958. National Technical Information Service, Washington, DC, USA, 2001. ix + 386 pp. URL https:/

                                        /nvlpubs.nist.gov/nistpubs/
                                        sp958-lide/cntsp958old.htm.

NIST Special Publication.

Leo:1980:SFR

[LLS80] Mario Leo, Rosario Antonio Leo, and Giulio Soliani. On a special function related to a class of solutions of certain nonlinear wave equations. Rend. Sem. Mat. Univ. Politec. Torino, 38(1):87– 110, 1980. ISSN 0373-1243.

Lorch:1973:SMP

[LMS73] Lee Lorch, M. E. Muldoon, and Peter Szego. Some monotonicity properties of Bessel functions. SIAM Journal on Mathematical Analysis, 4(2):385–392, May 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Ludwig:1970:SAM

[LO70] Donald Ludwig and Frank J. Olver, editors. Special Functions and Wave Propagation; a Collection of Papers Presented at the Symposia on Special functions and Wave Propagation, … at the 1969 SIAM national meeting held in Washington, DC, volume 6 of Studies in applied mathematics. SIAM Press, Philadelphia, PA, USA, 1970. LCCN ????

Lozano:1978:AAP

[LO78] C. Lozano and F. W. J. Olver.

Asymptotic approximations for parabolic cylinder functions. Journal of Physics B: Atomic and Molecular Physics, 11(18):L531– L533, September 1978. CO-DEN JPAMA4. ISSN 0953-4075 (print), 1361-6455 (electronic).

Lozier:1990:CPL

[LO90] Daniel W. Lozier and F. W. J.

Olver. Closure and precision in level-index arithmetic. SIAM Journal on Numerical Analysis, 27(5):1295–1304, 1990. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Lozier:1993:ABF

[LO93] Daniel W. Lozier and F. W. J.

Olver. Airy and Bessel functions by parallel integration of ODEs. In Sincovec et al. [SKL93], pages 530–538. ISBN 0-89871-315-3a. LCCN QA76.58 .S55 1993 v.1-2. Two volumes.

Lozier:1994:NESa

[LO94a] D. W. Lozier and F. W. J. Olver.

Numerical evaluation of special functions. Report NISTIR 5383, Computing and Applied Mathematics Laboratory, U. S. Department of Commerce, Washington, DC, USA, March 1994. 47

pp. URL https://math.nist.

gov/~DLozier/publications/
nistir5383.pdf.

Lozier:1994:NESb

[LO94b] D. W. Lozier and F. W. J. Olver. Numerical evaluation of special functions. In Gautschi [Gau94], pages 79–125. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993.

URL http://math.nist.gov/
mcsd/Reports/2001/nesf/. See

also SIAM Review, September 1995, 37(3), p. 483.

Liu:2014:BSI

[LOQZ14] Yue Liu, Peter J. Olver, Changzheng Qu, and Shuanghu Zhang. On the blow-up of solutions to the integrable modified Camassa–Holm equation. Analysis and Applications (Singapore), 12(4):355–368, June 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part I).

Leubner:1986:NUA

[LR86] C. Leubner and H. Ritsch. A note on the uniform asymptotic expansion of integrals with coalescing endpoint and saddle points. Journal of Physics A (Mathematical and General), 19(3):329– 335, 1986. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic). URL http:

//stacks.iop.org/0305-4470/
19/329.

Ludwig:1983:PA

[Lud83] Donald Ludwig. Parsimonious asymptotics. SIAM Journal on Applied Mathematics, 43(4): 664–672, August 1983. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712x (electronic).

Lin:2014:GAS

[LW14] Y. Lin and R. Wong. Global asymptotics of the Szegő–Askey polynomials. Analysis and Applications (Singapore), 12(6):727– 746, October 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part III).

MacGillivray:1975:UCP

[Mac75] A. D. MacGillivray. On Ursell’s combinatorial problem. Journal of Mathematical Physics, 16(9): 1735–1738, 1975. CODEN JMA-PAQ. ISSN 0022-2488.

McCann:1977:LBZ

[McC77] Roger C. McCann. Lower bounds for the zeros of Bessel functions. Proceedings of the American Mathematical Society, 64(1): 101–103, 1977. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Meyer:1980:EA

[Mey80] R. E. Meyer. Exponential asymptotics. SIAM Review, 22(2): 213–224, April 1980. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Meyer:1990:OTS

[Mey90] R. E. Meyer. Observable tunneling in several dimensions. In Wong [Won90], pages 299–327. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Miller:1952:MDC

[Mil52] J. C. P. Miller. A method for the determination of converging factors, applied to the asymptotic expansions for the parabolic cylinder functions. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 48:243–254, April 1952. CODEN PCPSA4. ISSN 0008-1981.

Muldoon:1979:ZCP

[Mul79] M. E. Muldoon. On the zeros of a cross-product of Bessel functions of different orders. Zeitschrift für Angewandte Mathematik und Mechanik, 59(6):272–273, 1979. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Muldoon:1972:SIW

[Mul73] M. E. Muldoon. Singular integrals whose kernels involve certain Sturm–Liouville functions. II. Indiana University Mathematics Journal, 22(1):51–63, 1972/73. CODEN IUMJAB. ISSN 0022-2518.

Mainardi:1990:AMS

[MV90] Francesco Mainardi and Giuliano Vitali. Applications of the method of steepest descents in wave-propagation problems. In Wong [Won90], pages 639– 651. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

McClure:1983:EBM

[MW83] J. P. McClure and R. Wong. Error bounds for multidimensional Laplace approximation. Journal of Approximation Theory, 37(4): 372–390, 1983. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Martin:1992:WAP

[MW92] P. A. Martin and G. R. Wickham, editors. Wave Asymptotics: the proceedings of the meeting to mark the retirement of Professor Fritz Ursell from the Beyer Chair of Applied Mathematics in the University of Manchester. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-41414-8 (hardcover). LCCN TC172 .W38 1992.

Nemes:2014:RPL

[Nem14] Gerg Nemes. The resurgence properties of the large-order asymptotics of the Hankel and Bessel functions. Analysis and Applications (Singapore), 12(4): 403–462, June 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part I).

OldeDaalhuis:1995:SPM

[OCK^{+}95] A. B. Olde Daalhuis, S. J. Chapman, J. R. King, J. R. Ockendon, and R. H. Tew. Stokes phenomenon and matched asymptotic expansions. SIAM Journal on Applied Mathematics, 55(6): 1469–1483, December 1995. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/26176.

Ojha:1973:SRG

[Ojh73] S. N. Ojha. Some results of generalised Legendre polynomials. Math. Education, 7:A84– A89, 1973.

Olver:2010:NHM

[OLBC10] Frank W. J. Olver, Daniel W. Lozier, Ronald F. Boisvert, and Charles W. Clark, editors. NIST Handbook of Mathematical Functions. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19225-0. xv + 951 pp. LCCN QA331 .N57 2010. US$99.00.

                                        URL http://dlmf.nist.gov/;
                                         http://www.cambridge.org/
                                        9780521140638.

Oliver:1968:EOE

[Oli68] J. Oliver. An extension of Olver’s error estimation technique for linear recurrence relations. Numerische Mathematik, 12(5):459– 467, December 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Olver:1946:NPH

[Olv46] F. W. J. Olver. Note on a paper of H. Bateman. Journal of Applied Physics, 17(12): 1127, December 1946. CO-DEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. Correction of numerical errors in [Bat46].

Olver:1949:TCS

[Olv49] F. W. J. Olver. Transformation of certain series occurring in aerodynamic interference calculations. Quarterly Journal of Mechanics and Applied Mathematics, 2:452–457, 1949. CO-DEN QJMMAV. ISSN 0033-5614 (print), 1464-3855 (electronic).

Olver:1950:NME

[Olv50] F. W. J. Olver. A new method for the evaluation of zeros of Bessel functions and of other solutions of second-order differential equations. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 46(4):570–580, 1950. CO-DEN PCPSA4. ISSN 0008-1981.

Olver:1951:FME

[Olv51] F. W. J. Olver. A further method for the evaluation of zeros of Bessel functions and some new asymptotic expansions for zeros of functions of large order. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 47(4):699– 712, October 1951. CODEN PCPSA4. ISSN 0008-1981.

Olver:1952:EZH

[Olv52a] F. W. J. Olver. The evaluation of zeros of high-degree polynomials. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 244(885):385– 415, April 1952. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Olver:1952:SNA

[Olv52b] F. W. J. Olver. Some new asymptotic expansions for Bessel functions of large orders. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 48(3):414–427, July 1952. CODEN PCPSA4. ISSN 0008-1981.

Olver:1953:NAE

[Olv53] F. W. J. Olver. Note on the asymptotic expansion of generalized hypergeometric functions. Journal of the London Mathematical Society, 28(4):462–464, October 1953. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Olver:1954:AEB

[Olv54a] F. W. J. Olver. The asymptotic expansion of Bessel functions of large order. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 247: 328–368, 1954. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Olver:1954:AEL

[Olv54b] F. W. J. Olver. The asymptotic expansion of Legendre functions of large order. Abstract. In ????, editor, Proceedings of the International Congress of Mathematicians, volume 2, page 152. North-Holland Publishing Co., Amsterdam, The Netherlands, 1954.

Olver:1954:ASL

[Olv54c] F. W. J. Olver. The asymptotic solution of linear differential equations of the second order for large values of a parameter. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 247:307–327, 1954. CODEN PTRMAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Olver:1956:ASL

[Olv56a] F. W. J. Olver. The asymptotic solution of linear differential equations of the second order in a domain containing one transition point. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 249:65–97, 1956. CODEN PTRMAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Olver:1956:EAS

[Olv56b] F. W. J. Olver. Errors in asymptotic solutions of linear ordinary differential equations. Quarterly of Applied Mathematics, 14:218– 219, 1956. CODEN QAMAAY.

ISSN 0033-569X (print), 1552-4485 (electronic).

Olver:1958:UAE

[Olv58]  F. W. J. Olver.  Uniform

asymptotic expansions of solutions of linear second-order differential equations for large values of a parameter. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 250: 479–517, 1958. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Olver:1959:LDE

[Olv59a] F. W. J. Olver. Linear differential equations of the second order with a large parameter. Journal of the Society for Industrial and Applied Mathematics, 7(3): 306–310, September 1959. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Olver:1959:UAE

[Olv59b] F. W. J. Olver. Uniform asymptotic expansions for Weber parabolic cylinder functions of large orders. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 63B:131–169, 1959. CODEN JNBBAU. ISSN 0022-4340.

Olver:1960:BFP

[Olv60] F. W. J. Olver, editor. Bessel functions: (Part 3: Zeros and associated values), volume 7 of Royal Society Mathematical Tables. Cambridge University Press, Cambridge, UK, 1960. lx + 79 pp. LCCN QA351 .H3.

Olver:1961:EBL

[Olv61a] F. W. J. Olver. Error bounds for the Liouville–Green (or WKB) approximation. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 57(4):790–810, October 1961. CODEN PCPSA4. ISSN 0008-1981.

Olver:1961:TIP

[Olv61b] F. W. J. Olver. Two inequalities for parabolic cylinder functions. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 57(4): 811–822, October 1961. CODEN PCPSA4. ISSN 0008-1981.

Olver:1961:U

[Olv61c] Frank William John Olver.

[unknown]. D.Sc. dissertation, Queen Mary University of London, London, UK,

1961. URL https://discover.
libraryhub.jisc.ac.uk/search?
author=Olver%2C%20F&date=1961&
rn=1.

Olver:1962:TBF

[Olv62] F. W. J. Olver. Tables for Bessel Functions of Moderate or Large Orders, volume 6 of National Physical Laboratory Mathematical Tables. Her Majesty’s Stationery Office, London, UK, 1962. iii + 51 pp.

Olver:1963:EBF

[Olv63a] F. W. J. Olver. Error bounds for first approximations in turningpoint problems. Journal of the Society for Industrial and Applied Mathematics, 11(3):748– 772, September 1963. CO-DEN JSIMAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Olver:1963:RPS

[Olv63b] F. W. J. Olver. Recent publications: A Survey of Numerical Analysis. American Mathematical Monthly, 70(2):224–225, February 1963. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Olver:1964:BFI

[Olv64a] F. W. J. Olver. Bessel functions of integer order. In Abramowitz and Stegun [AS64], pages 355– 434. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://www.

convertit.com/Go/ConvertIt/
Reference/AMS55.ASP in bitmap

image format.

Olver:1964:EAM

[Olv64b] F. W. J. Olver. Error analysis of Miller’s recurrence algorithm. Mathematics of Computation, 18 (85):65–74, 1964. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olver:1964:EBAb

[Olv64c] F. W. J. Olver. Error bounds for asymptotic expansions in turning-point problems. Journal of the Society for Industrial and Applied Mathematics, 12(1):200– 214, March 1964. CODEN JSI-MAV. ISSN 0368-4245 (print), 2168-3484 (electronic).

Olver:1964:EBAa

[Olv64d] F. W. J. Olver. Error bounds for asymptotic expansions, with an application to cylinder functions of large argument. In Wilcox [Wil64], pages 163–183. Publication No. 13 of the Mathematics Research Center, United States Army, The University of Wisconsin. Dedicated to Professor Rudolph E. Langer.

Olver:1965:BRR

[Olv65a] F. W. J. Olver. Book review: A.

R. Curtis, Coulomb Wave Functions, Royal Society Mathematical Tables, Volume 11, Cambridge University Press, New York, 1964, xxxv + 209 pp. Mathematics of Computation, 19 (90):341–342, April 1965. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.ams.
org/journals/mcom/1965-19-
090/S0025-5718-65-99950-3/
S0025-5718-65-99950-3.pdf.

Olver:1965:EAPb

[Olv65b] F. W. J. Olver. Error analysis of phase-integral methods. I. General theory for simple turning points. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 69B:271–290, 1965. CODEN JNBBAU. ISSN 0022-4340.

Olver:1965:EAPa

[Olv65c] F. W. J. Olver. Error analysis of phase-integral methods. II. Application to wave-penetration problems. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 69B:291–300, 1965. CODEN JNBBAU. ISSN 0022-4340.

Olver:1965:EBAa

[Olv65d] F. W. J. Olver. Error bounds for asymptotic expansions of special functions in the complex plane. In Rall [Ral65], pages 55–75. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5–7, 1964.

Olver:1965:ASS

[Olv65e] F. W. J. Olver. On the asymptotic solution of second-order differential equations having an irregular singularity of rank one, with an application to Whittaker functions. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):225–243, January 1965. ISSN 0887-459X (print), 2168-3581 (electronic).

Olver:1966:ESF

[Olv66] F. W. J. Olver. Evaluation of special functions. In ????, editor, Proceedings of IFIP Congress 65, volume 2, pages 604–605. Spartan Books, Washington, DC, USA,

Olver:1967:BRA

[Olv67a] F. W. J. Olver. Book review: Asymptotic Expansions, by H. A. Lauwerier: Mathematisch Centrum, Amsterdam, 1966. Quarterly of Applied Mathematics, 25(1):117, 1967. CODEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic). URL https:/

/www.ams.org/journals/qam/
1967-25-01/qam-25-1-print-
matter.pdf.

Olver:1967:BSS

[Olv67b] F. W. J. Olver. Bounds for the solutions of second-order linear difference equations. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 71B:161–166, 1967. CO-DEN JNBBAU. ISSN 0022-4340.

Olver:1967:EMA

[Olv67c] F. W. J. Olver. An extension of Miller’s algorithm. Aplikace Matematiky, 13:174–176, 1967. CODEN APMTAK. ISSN 0373-6725.

Olver:1967:NSS

[Olv67d] F. W. J. Olver. Numerical solution of second-order linear difference equations. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 71B(2–3):111–129, April 1967. CODEN JNBBAU. ISSN 0022-

4340. URL https://nvlpubs.
nist.gov/nistpubs/jres/71B/
jresv71Bn2-3p111_A1b.pdf.

Olver:1968:EBL

[Olv68] F. W. J. Olver. Error bounds for the Laplace approximation for definite integrals. Journal of Approximation Theory, 1:293– 313, 1968. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Olver:1970:PA

[Olv70a] F. W. J. Olver. A paradox in asymptotics. SIAM Journal on Mathematical Analysis, 1(4): 533–534, November 1970. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1970:WSDa

[Olv70b] F. W. J. Olver. Why steepest descents? SIAM Review, 12:228– 247, 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1970:WSDb

[Olv70c] F. W. J. Olver. Why steepest descents? In Ludwig and Olver [LO70], pages 44–63. LCCN ????

Olver:1972:BRS

[Olv72] Frank W. J. Olver. Book review: The Special Functions and Their Approximations (Vols. I & II Yudell L. Luke). SIAM Review, 14(1):181–183, ???? 1972. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1974:AM

[Olv74a] F. W. J. Olver. Asymptotic methods. In C. E. Pearson, editor, Handbook of Applied Mathematics, pages 645–710. Van Nostrand Reinhold, New York, NY, USA, 1974.

Olver:1974:ASF

[Olv74b] F. W. J. Olver. Asymptotics and Special Functions. Computer Science and Applied Mathematics. Academic Press, New York, USA, 1974. ISBN 0-12-525850-X. xvi + 572 pp. LCCN QA351 .O481 1974.

Olver:1974:EBS

[Olv74c] F. W. J. Olver. Error bounds for stationary phase approximations. SIAM Journal on Mathematical Analysis, 5(1):19–29, February 1974. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1975:LFB

[Olv75a] F. W. J. Olver. Legendre functions with both parameters large. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 278:175–185, 1975. CODEN PTRMAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Olver:1975:REA

[Olv75b] F. W. J. Olver. Recent error analysis of asymptotic solutions of linear differential equations. In Henry Albert Antosiewicz, editor, International Conference on Differential Equations: proceedings of an international conference on differential equations held at the University of Southern California [Los Angeles], September 3–7, 1974, pages 636–645. Academic Press, New York, USA, 1975. ISBN 0-12-059650-4, 1-4832-5913-7 (ebook). LCCN QA371 .I54 1974.

Olver:1975:SOL

[Olv75c] F. W. J. Olver. Second-order linear differential equations with two turning points. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 278: 137–174, 1975. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Olver:1975:UPA

[Olv75d] F. W. J. Olver. Unsolved problems in the asymptotic estimation of special functions. In Theory and application of special functions (Proceedings of an Advanced Seminar Sponsored by the Mathematics Research Center, the University of Wisconsin– Madison, March 31–April 2, 1975), pages 99–142. Academic Press, New York, USA, 1975.

Olver:1976:IEB

[Olv76a] F. W. J. Olver. Improved error bounds for second-order differential equations with two turning points. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 80B(4):437–440, 1976. CODEN JNBBAU. ISSN 0022-4340.

Olver:1976:UAE

[Olv76b] F. W. J. Olver. Uniform asymptotic expansions and singular perturbations. In R. E. O’Malley, Jr., editor, Asymptotic methods and singular perturbations (Proceedings of the SIAM-AMS Symposium on Applied Mathematics, New York, 1976), volume X of SIAM-AMS Proc., pages 105– 117. American Mathematical Society, Providence, RI, USA, 1976.

Olver:1977:CFSa

[Olv77a] F. W. J. Olver. Connection formulas for second-order differential equations having an arbitrary number of turning points of arbitrary multiplicities. SIAM Journal on Mathematical Analysis, 8 (4):673–700, 1977. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1977:CFSb

[Olv77b] F. W. J. Olver. Connection formulas for second-order differential equations with multiple turning points. SIAM Journal on Mathematical Analysis, 8(1):127– 154, 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1977:SOD

[Olv77c] F. W. J. Olver. Second-order differential equations with fractional transition points. Transactions of the American Mathematical Society, 226:227–241, 1977. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Olver:1978:CVV

[Olv78a] F. Olver. Vvedenie v asimptoticheskie metody i spetsial′nye funktsii (Russian) [Introduction to asymptotic methods and special functions]. “Nauka”, Moscow, USSR, 1978. 375 pp. Translation by Ju. A. Bryčkov of [Olv74b]. Edited by A. P. Prudnikov.

Olver:1978:GCF

[Olv78b] F. W. J. Olver. General connection formulae for Liouville–Green approximations in the complex plane. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 289(1364):501– 548, 1978. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic). See corrigendum [Olv79].

Olver:1978:NAE

[Olv78c] F. W. J. Olver. A new approach to error arithmetic. SIAM Journal on Numerical Analysis, 15 (2):368–393, 1978. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Olver:1978:SCA

[Olv78d] F. W. J. Olver. Sufficient conditions for Ackerberg–O’Malley resonance. SIAM Journal on Mathematical Analysis, 9(2):328– 355, 1978. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1979:C

[Olv79] Frank W. J. Olver. Corrigendum. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 290(1377):686, February 1979. CODEN PTR-MAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic). See [Olv78b].

Olver:1980:AAE

[Olv80a] F. W. J. Olver. Asymptotic approximations and error bounds. SIAM Review, 22(2):188–203, 1980. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1980:UAG

[Olv80b] F. W. J. Olver. Unrestricted algorithms for generating elementary functions. In G. Alefeld and R. D. Grigorieff, editors, Fundamentals of Numerical Computation (Computer-Oriented Numerical Analysis) (Proceedings of the Conference at the Technical University of Berlin, Berlin, 1979), volume 2 of Comput. Suppl., pages 131–140. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 3-211-81566-X.

Olver:1980:WFB

[Olv80c] F. W. J. Olver. Whittaker functions with both parameters large: uniform approximations in terms of parabolic cylinder functions. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and Physical Sciences, 86(3–4):213–234, 1980. CO-DEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic).

Olver:1980:GCF

[Olv80d] Frank W. J. Olver. The general connection-formula problem for linear differential equations of the second order. In R. E. Meyer and S. V. Parter, editors, Singular perturbations and asymptotics (Proceedings of an Advanced Seminar, Mathematics Research Center, University of Wisconsin, Madison, Wisconsin, 1980), volume 45 of Publication of the Mathematics Research Center, University of Wisconsin, pages 317–343. Academic Press, New York, USA, 1980. ISBN 0-12-493260-6.

Olver:1982:FDR

[Olv82] F. W. J. Olver. Further developments of Rp and Ap error analysis. IMA Journal of Numerical Analysis, 2(3):249–274, July 1982. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Olver:1983:AM

[Olv83a] F. W. J. Olver. Asymptotic methods. In C. E. Pearson, editor, Handbook of Applied Mathematics, pages 631–696. Van Nostrand Reinhold, New York, NY, USA, second edition, 1983.

Olver:1983:EAC

[Olv83b] F. W. J. Olver. Error analysis of complex arithmetic. In H. Werner, L. Wuytack, E. Ng, and H. J. Bunger, editors, Computational aspects of complex analysis (Braunlage, 1982), volume 102 of NATO Advanced Science Institute Series C: Mathematical and Physical Sciences, pages 279–292. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1571-8.

Olver:1983:EBA

[Olv83c] F. W. J. Olver. Error bounds for arithmetic operations on computers without guard digits. IMA Journal of Numerical Analysis, 3 (2):153–160, 1983. CODEN IJ-NADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Olver:1984:BRM

[Olv84] F. W. J. Olver. Book review:

Mathematics for the Analysis of Algorithms, by D. H. Greene and D. E. Knuth: Progress in Computer Science, v. 1 (second edition), Birkhäuser, Boston, 1982. COMPEL, 3(??):53–54, ???? 1984.

Olver:1985:UAM

[Olv85] F. W. J. Olver. Unrestricted algorithms for mathematical functions. Rend. Sem. Mat. Univ. Politec. Torino, ??(??):237–248, ???? 1985. ISSN 0373-1243. International conference on special functions: theory and computation (Turin, 1984).

Olver:1986:BRC

[Olv86a] F. W. J. Olver. Book review:

Computation with Recurrence Relations, by Jet Wimp: Pitman, Boston, 1984. Mathematics of Computation, 47(175):371– 372, 1986. CODEN MCMPAF.

ISSN 0025-5718 (print), 1088-6842 (electronic). URL https:/

                                        /www.ams.org/journals/mcom/
                                        1986-47-175/S0025-5718-86-
                                        99764-4.

Olver:1986:EBP

[Olv86b] F. W. J. Olver. Error bounds for polynomial evaluation and complex arithmetic. IMA Journal of Numerical Analysis, 6(3):373– 379, 1986. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Olver:1986:LIN

[Olv86c] F. W. J. Olver. The level-index number system. Bulletin of the Institute of Mathematics and its Applications, 22(11–12):182–185, 1986. CODEN IMTABW. ISSN 0905-5628.

Olver:1986:SEB

[Olv86d] F. W. J. Olver. Simplified error bounds for Newton’s rule. Universidad del Zulia. Facultad de Ingeniería. Revista Técnica, 9(2): 71–77, 1986. ISSN 0254-0770.

Olver:1986:BRL

[Olv86e] Frank W. J. Olver. Book review: Linear Turning Point Theory (Wolfgang Wasow). SIAM Review, 28(3):433–435, ???? 1986. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1987:CCA

[Olv87] F. W. J. Olver. A closed computer arithmetic. In Irwin and Stefanelli [IS87], pages 139–143. ISBN 0-8186-0774-2 (paperback), 0-8186-4774-4 (microfiche), 0-8186-8774-6 (case). LCCN QA 76.9 C62 S95

1987. URL http://www.acsel-
lab.com/arithmetic/arith8/
papers/ARITH8_Olver.pdf.

Olver:1988:EBL

[Olv88] F. W. J. Olver. Error bounds for linear recurrence relations. Mathematics of Computation, 50 (182):481–499, April 1988. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olver:1989:BRT

[Olv89a] F. W. J. Olver. Book review:

Theory of Difference Equations: Numerical Methods and Applications, by V. Lakshmikantham and D. Trigiante: Mathematics in Science and Engineering, v. 181, Academic Press, New York, 1988. American Scientist, 77(3): 294–295, May/June 1989. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL https://www.jstor.org/
stable/27855786.

Olver:1989:CP

[Olv89b] F. W. J. Olver. Closure and precision. In Turner [Tur89], pages 124–130. ISBN 3-540-51645-X. Lectures from the Third S.E.R.C. Numerical Analysis Summer School held at the University of Lancaster, Lancaster, 1987.

Olver:1989:GEL

[Olv89c] F. W. J. Olver. Generalized exponentials and logarithms; surface fitting; conclusions. In Turner [Tur89], pages 156–164. ISBN 3-540-51645-X. Lectures from the Third S.E.R.C. Numerical Analysis Summer School held at the University of Lancaster, Lancaster, 1987.

Olver:1990:CAS

[Olv90a] F. Olver. Asimptotika i spetsial′nye funktsii. (Russian) [Asymptotics and special functions]. “Nauka”, Moscow, USSR, 1990. ISBN 5-02-014228-X. 528 pp. Translation by Yu. A. Brychkov of [Olv74b]. Translation edited and with a preface by A. P. Prudnikov.

Olver:1990:REA

[Olv90b] F. W. J. Olver. Rounding errors in algebraic processes—in levelindex arithmetic. In Cox and Hammarling [CH90], pages 197– 205. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th– 10th July 1987.

Olver:1990:SPC

[Olv90c] Frank W. J. Olver. On Stokes’ phenomenon and converging factors. In Wong [Won90], pages 329–356. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Olver:1991:RUB

[Olv91a] F. W. J. Olver. Book review: Mathematical Function Library for Microsoft-FORTRAN, by United Laboratories, Inc.: Wiley, New York, 1989. Mathematics of Computation, 56(194): 879–885, April 1991. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL https://www.
ams.org/journals/mcom/1991-
56-194/S0025-5718-91-99751-
6.

Olver:1991:UEIa

[Olv91b] F. W. J. Olver. Uniform, exponentially improved, asymptotic expansions for the confluent hypergeometric function and other integral transforms. SIAM Journal on Mathematical Analysis, 22 (5):1475–1489, 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1991:UEIb

[Olv91c] F. W. J. Olver. Uniform, exponentially improved, asymptotic expansions for the generalized exponential integral. SIAM Journal on Mathematical Analysis, 22 (5):1460–1474, 1991. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1992:CF

[Olv92] F. W. J. Olver. Converging factors. In Martin and Wickham [MW92], pages 54–68. ISBN 0-521-41414-8 (hardcover). LCCN TC172 .W38 1992.

Olver:1993:EIAa

[Olv93a] F. W. J. Olver. Exponentially improved asymptotic solutions of ordinary differential equations. In B. D. Sleeman and R. J. Jarvis, editors, Ordinary and partial differential equations, Vol. IV (Dundee, 1992), volume 289 of Pitman Res. Notes Math. Ser., pages 238–247. Longman Scientific and Technical, Harlow, Essex, 1993. ISBN 0-582-09137-3.

Olver:1993:EIAb

[Olv93b] F. W. J. Olver. Exponentiallyimproved asymptotic solutions of ordinary differential equations. I. The confluent hypergeometric function. SIAM Journal on Mathematical Analysis, 24(3): 756–767, May 1993. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1994:AEC

[Olv94a] F. W. J. Olver. Asymptotic expansions of the coefficients in asymptotic series solutions of linear differential equations. Methods and Applications of Analysis, 1(1):1–13, 1994. ISSN 1073-2772 (print), 1945-0001 (electronic).

Olver:1994:GEI

[Olv94b] F. W. J. Olver. The generalized exponential integral. In Approximation and computation (West Lafayette, IN, 1993), volume 119 of Internat. Ser. Numer.

Math., pages 497–510. Birkhäuser Boston Inc., Cambridge, MA, USA, 1994. ISBN 0-8176-3753-2.

Olver:1995:AER

[Olv95] F. W. J. Olver. On an asymptotic expansion of a ratio of gamma functions. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 95(1):5–9, 1995. CODEN PRIAAK. ISSN 0035-8975.

Olver:1997:ASL

[Olv97a] F. W. J. Olver. Asymptotic solutions of linear ordinary differential equations at an irregular singularity of rank unity. Methods and Applications of Analysis, 4(4):375–403, 1997. ISSN 1073-2772 (print), 1945-0001 (electronic).

Olver:1997:BRS

[Olv97b] Frank W. J. Olver. Book review: Special Functions: An Introduction to the Classical Functions of Mathematical Physics, by Nico M. Temme. SIAM Review, 39(2):355–356, June 1997. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olver:1998:AAN

[Olv98a] F. W. J. Olver. Advances in the asymptotic and numerical solution of linear ordinary differential equations. In H-H. Dai and P. L. Sachdev, editors, Recent advances in differential equations (Kunming, 1997), volume 386 of Pitman Res. Notes Math. Ser., pages 3–14. Longman Scientific and Technical, Harlow, Essex, 1998. ISBN 0-582-32219-7.

Olver:1998:PPS

[Olv98b] Frank William John Olver. Practical problems in the solution of high-order linear differential equations. In Tricomi’s ideas and contemporary applied mathematics (Rome, 28–29 November, 1997/Turin, 1–2 December 1997), volume 147 of Atti Convegni Lincei, pages 239–249. Accademia Nazionale dei Lincei, Rome, Italy, 1998. ISSN 0391-805X. LCCN QA299.6 .T75 1998.

Olver:1999:UAS

[Olv99]  F. W. J. Olver.  On the

uniqueness of asymptotic solutions of linear differential equations. Methods and Applications of Analysis, 6(2):165–174, 1999. ISSN 1073-2772 (print), 1945-0001 (electronic). Dedicated to Richard A. Askey on the occasion of his 65th birthday, Part II.

Olver:2000:RDE

[Olv00] F. W. J. Olver. Resurgence in difference equations, with an application to Legendre functions. In Dunkl et al. [DIW00], pages 221– 235. ISBN 981-02-4393-6. LCCN QA351.P76 2000; QC20.7.F87 P76 2000.

Olver:2004:CCM

[Olv04] Frank Olver. Charles Clenshaw: Mathematician with a flair for numerical analysis. The Guardian, ??(??):??, October 28, 2004. ISSN 0261-3077 (print), 1756-3224 (elec-

tronic). URL https://www.
theguardian.com/news/2004/
oct/29/guardianobituaries.
obituaries.

Olver:2009:HP

[Olv09] Frank W. J. Olver. Home page.

Archived Web site, July 13, 2009.

URL https://web.archive.
org/web/20100530002944/http:
//www.ipst.umd.edu/researchandfaculty/
Olver/.

Olver:2010:ARF

[Olv10] F. W. J. Olver. Airy and related functions. In NIST handbook of mathematical functions, pages 193–213. U.S. Department of Commerce, Washington, DC, USA, 2010. ISBN 0-521-14063-3.

Olver:2014:MIM

[Olv14a] F. W. J. Olver. Mathematics that has intrigued me. Analysis and Applications (Singapore), 12(4): 341–354, 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Lecture delivered at Asymptotics and Applied Analysis, Conference in Honor of Frank W. J. Olver’s 75th Birthday, January 10–14, 2000, San Diego State University, San Diego, California.

Olver:2014:CVF

[Olv14b] Sheehan Olver. Change of variable formulas for regularizing slowly decaying and oscillatory Cauchy and Hilbert transforms. Analysis and Applications (Singapore), 12(4):369–384, June 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part I).

Olver:1990:CCA

[Olv09] F. W. J. Olver. A closed computer arithmetic. In Computer Arithmetic [Swa90], pages 360– 364. ISBN 0-8186-8945-5. LCCN QA76.9 .C62C66 1990. This is part of a two-volume collection of influential papers on the design of computer arithmetic.

Olver:2010:BF

[OM10] F. W. J. Olver and L. C. Maximon. Bessel functions. In NIST handbook of mathematical functions, pages 215–286. U.S. Department of Commerce, Washington, DC, USA, 2010. ISBN 0-521-14063-3.

OldeDaalhuis:1994:EIA

[OO94] A. B. Olde Daalhuis and F. W. J. Olver. Exponentially improved asymptotic solutions of ordinary differential equations. II. Irregular singularities of rank one. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 445(1923): 39–56, 1994. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

OldeDaalhuis:1995:HSS

[OO95a] A. B. Olde Daalhuis and F. W. J. Olver. Hyperasymptotic solutions of second-order linear differential equations. I. Methods and Applications of Analysis, 2(2): 173–197, 1995. ISSN 1073-2772 (print), 1945-0001 (electronic).

OldeDaalhuis:1995:CSM

[OO95b] A. B. Olde Daalhuis and F. W. J.

Olver. On the calculation of Stokes multipliers for linear differential equations of the second order. Methods and Applications of Analysis, 2(3):348–367, 1995. ISSN 1073-2772 (print), 1945-0001 (electronic).

OldeDaalhuis:1998:ANS

[OO98] A. B. Olde Daalhuis and F. W. J.

Olver. On the asymptotic and numerical solution of linear ordinary differential equations. SIAM Review, 40(3):463– 495, 1998. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Olkha:1971:GBF

[OR71] G. S. Olkha and P. N. Rathie.

On a generalized Bessel function and an integral transform. Mathematische Nachrichten, 51(1–6): 231–240, 1971. CODEN MTM-NAQ. ISSN 0022-2526 (print), 1467-9590 (electronic).

Olver:1965:EBAb

[OS65] F. W. J. Olver and F. Stenger.

Error bounds for asymptotic solutions of second-order differential equations having an irregular singularity of arbitrary rank. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 2(2):244–249, ???? 1965. CODEN ???? ISSN 0887-459X (print), 2168-3581 (electronic).

URL http://links.jstor.org/

                                        sici?sici=0887-459X(1965)2:
                                        2<244:EBFASO>2.0.CO%3B2-N.

Olver:1972:NBR

[OS72] F. W. J. Olver and D. J. Sookne. Note on backward recurrence algorithms. Mathematics of Computation, 26(120):941–947, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olver:1983:ALF

[OS83] F. W. J. Olver and J. M. Smith. Associated Legendre functions on the cut. Journal of Computational Physics, 51(3):502–518, September 1983. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Olver:1987:ILI

[OT87] F. W. J. Olver and P. R. Turner. Implementation of level-index arithmetic using partial table look-up. In Irwin and Stefanelli [IS87], pages 144–147. ISBN 0-8186-0774-2 (paperback), 0-8186-4774-4 (microfiche), 0-8186-8774-6 (case). LCCN QA 76.9 C62 S95

                                        1987. URL http://www.acsel-
                                        lab.com/arithmetic/arith8/
                                        papers/ARITH8_Olver_Turner.
                                        pdf.

Olver:1982:PEB

[OW82] F. W. J. Olver and J. H. Wilkinson. A posteriori error bounds for Gaussian elimination. IMA Journal of Numerical Analysis, 2 (4):377–406, 1982. CODEN IJ-NADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Olver:2010:AA

[OW10] F. W. J. Olver and R. Wong.

Asymptotic approximations. In NIST handbook of mathematical functions, pages 41–70. U.S. Department of Commerce, Washington, DC, USA, 2010. ISBN 0-521-14063-3.

Parodi:1974:SDT

[Par74] Maurice Parodi. Sur la série dont le terme général est égal à la somme des puissances μ-ième des zéros d’un polynôme de Legendre. (French) [On the series whose general term is equal to the sum of the μ-th powers of the zeros of a Legendre polynomial]. C. R. Acad. Sci. Paris Sér. A, 278: 153–154, 1974. ISSN 0302-8429.

Paris:1975:ABS

[Par75] R. B. Paris. The asymptotic behaviour of solutions of the differential equation d^{4}u/dz^{4} + [z^{2}(d^{2}u/dz^{2}) + az(du/dz) + bu] = 0. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 346(1645):171–207, 1975. CO-DEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Paris:1990:API

[Par90]  R. B. Paris.  The asymp-

totics of Pearcey’s integral for complex variables. In Wong [Won90], pages 653–667. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Paris:1992:SSP

[Par92] R. B. Paris. Smoothing of the Stokes phenomenon for highorder differential equations. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 436(1896):165–186, 1992. CO-DEN PRLAAZ. ISSN 0962-8444 (print), 2053-9177 (electronic).

Paris:1994:GPI

[Par94] R. B. Paris. A generalization of Pearcey’s integral. SIAM Journal on Mathematical Analysis, 25 (2):630–645, March 1994. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/23054.

Painter:1981:NCM

[PM81] J. F. Painter and R. E. Meyer. New connection method across more general turning points. Bulletin of the American Mathematical Society (new series), 4 (3):335–338, 1981. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Pryce:1984:NMR

[Pry84] J. D. Pryce. A new measure of relative error for vectors. SIAM Journal on Numerical Analysis, 21(1):202–215, February 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Paris:1980:AES

[PW80] R. B. Paris and A. D. Wood. The asymptotic expansion of solutions of the differential equation u^{iv} + λ^{2}[(z^{2} + c)u′′ + azu′ + bu] = 0 for large |z|. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 293(1404): 511–533, 1980. CODEN PTR-MAD, PTMSFB. ISSN 0080-4614 (print), 2054-0272 (electronic).

Qu:1988:SCL

[QW88] C. K. Qu and R. Wong. Szegő’s conjecture on Lebesgue constants for Legendre series. Pacific Journal of Mathematics, 135(1):157– 188, 1988. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http:

//projecteuclid.org/euclid.
pjm/1102688349.

Rall:1965:EDCb

[Ral65] L. B. Rall, editor. Error in Digital Computation, volume 2. Wiley, New York, NY, USA, 1965. Proceedings of an advanced seminar conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, October 5– 7, 1964.

Rudraiah:1990:AMM

[RCB90] N. Rudraiah, O. P. Chandna, and R. M. Barron. Asymptotic methods in magnetoconvection. In Wong [Won90], pages 669–683. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Reid:1990:ATO

[Rei90] W. H. Reid. On the asymptotic theory of the Orr–Sommerfeld problem. In Wong [Won90], pages 379–391. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Roy:2010:EF

[RO10] R. Roy and F. W. J. Olver. Elementary functions. In NIST handbook of mathematical functions, pages 103–134. U.S. Department of Commerce, Washington, DC, USA, 2010. ISBN 0-521-14063-3.

Roy:2010:AAM

[ROAW10] R. Roy, F. W. J. Olver, R. A. Askey, and R. Wong. Algebraic and analytic methods. In NIST handbook of mathematical functions, pages 1–39. U.S. Department of Commerce, Washington, DC, USA, 2010. ISBN 0-521-14063-3.

S:2015:NFS

[S.15] Cesar A. Gómez S. A nonlinear fractional Sharma–Tasso– Olver equation: New exact solutions. Applied Mathematics and Computation, 266(??):385– 389, September 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300315006980.

Salzer:1983:NDG

[Sal83] Herbert E. Salzer. Note on the Dočev–Grosswald asymptotic series for generalized Bessel polynomials. Journal of Computational and Applied Mathematics, 9(2):131–135, 1983. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). See errata [Sal84].

Salzer:1984:END

[Sal84] H. E. Salzer. Errata: “Note on the Dočev–Grosswald asymptotic series for generalized Bessel polynomials”. Journal of Computational and Applied Mathematics, 10(1):133, February 1984. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). See [Sal83].

Schmid:1969:BRF

[Sch69] E. Schmid. Book review: F.

L. Bauer and A. S. Householder and F. J. Olver and H. Rutishauser and K. Samelson and E. Stiefel, Handbook for Automatic computation, Vol. I. Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen mit besonderer Berücksichtigung der Anwendungsgebiete, Band 135. Computing, 4(3):276, September 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Stoyanov:1990:AEI

[SFB90] B. J. Stoyanov, R. A. Farrell, and J. F. Bird. Asymptotic expansions of integrals of two Bessel functions. In Wong [Won90], pages 723–740. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Sharapudinov:1988:AFH

[Sha88] I. I. Sharapudinov. An asymptotic formula having no remainder term for the orthogonal Hahn polynomials of discrete variable. Math. Balkanica (N.S.), 2(4): 314–318, 1988. ISSN 0205-3217.

Shih:1990:ILS

[Shi90] Shagi-Di Shih. The interior layer structure for a linear parabolic problem with discontinuous data. In Wong [Won90], pages 685– 696. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Shang:2011:BTA

[SHY11] Yadong Shang, Yong Huang, and Wenjun Yuan. Bäcklund transformations and abundant exact explicit solutions of the Sharma– Tasso–Olver equation. Applied Mathematics and Computation, 217(17):7172–7183, May 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300311001573.

Sibuya:1990:GPF

[Sib90] Yasutaka Sibuya. Gevrey property of formal solutions in a parameter. In Wong [Won90], pages 393–401. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Siegel:1990:AES

[Sie90] D. Siegel. Asymptotic expansions of solutions to Δu − u = f at infinity. In Wong [Won90], pages 697–705. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Sikkema:1983:ACO

[Sik83] P. C. Sikkema. On approximation with convolution operators containing three parameters. Delft progress report, 8(4):298– 310, 1983. CODEN DPRED2. ISSN 0304-985X.

Skinner:1980:UVE

[Ski80] L. A. Skinner. Uniformly valid expansions for Laplace integrals. SIAM Journal on Mathematical Analysis, 11(6):1058–1067, November 1980. CODEN SJ-MAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See erratum [Ski81].

Skinner:1981:EUV

[Ski81] L. A. Skinner. Erratum: “Uniformly valid expansions for Laplace integrals”. SIAM Journal on Mathematical Analysis, 12(3):487, May 1981. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). See [Ski80].

Skinner:1993:MAE

[Ski93] L. A. Skinner. Matched asymptotic expansions of integrals. IMA Journal of Applied Mathematics, 50(1):77–90, ???? 1993. CODEN IJAMDM. ISSN 0272-4960 (print), 1464-3634 (electronic).

Sincovec:1993:PSS

[SKL93] Richard F. Sincovec, David E. Keyes, and Michael R. Leuze, editors. Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, held March 22–24, 1993, in Norfolk, VA, USA. SIAM Press, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3a. LCCN QA76.58 .S55 1993 v.1-2. Two volumes.

Sadowski:1972:UOA

[SL72] W. L. Sadowski and Daniel W. Lozier. Use of Olver’s algorithm to evaluate certain definite integrals of plasma physics involving Chebyshev polynomials. Journal of Computational Physics, 10(3):607–613, December 1972. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/002199917290054X.

Smith:1986:LGA

[Smi86] Donald R. Smith. Liouville– Green approximations via the Riccati transformation. Journal of Mathematical Analysis and Applications, 116(1):147– 165, 1986. CODEN JMANAK.

ISSN 0022-247x (print), 1096-0813 (electronic).

Smith:1990:RAA

[Smi90]  Donald R. Smith.  A Ric-

cati approach to the Airy equation. In Wong [Won90], pages 403–415. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Smith:1981:ERA

[SOL81] J. M. Smith, F. W. J. Olver, and Daniel W. Lozier. Extendedrange arithmetic and normalized Legendre polynomials. ACM Transactions on Mathematical Software, 7(1):93–105, March 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Soni:1980:UAE

[Son80] Kusum Soni. On uniform asymptotic expansions of finite Laplace and Fourier integrals. Proceedings of the Royal Society of Edinburgh. Section A, Mathematical and Physical Sciences, 85 (3–4):299–305, 1980. CODEN PEAMDU. ISSN 0308-2105 (print), 1473-7124 (electronic).

Salchev:1973:DZC

[SP73] L. Z. Salchev and V. B. Popov.

Determination of the zeros of a cross-product Bessel function. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 74(3): 477–483, November 1973. CO-DEN PCPSA4. ISSN 0008-1981.

Shang:2008:AEE

[SQHY08] Yadong Shang, Jinghong Qin, Yong Huang, and Wenjun Yuan. Abundant exact and explicit solitary wave and periodic wave solutions to the Sharma–Tasso– Olver equation. Applied Mathematics and Computation, 202(2): 532–538, August 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0096300308001379.

Soni:1990:SPA

[SS90] K. Soni and R. P. Soni. A system of polynomials associated with the Chester, Friedman, and Ursell technique. In Wong [Won90], pages 417–440. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0647/90002810-
                                        d.html.

Stadje:1983:PPS

[Sta83] Wolfgang Stadje. Probabilistic proofs of some formulas for Bessel functions. Nederl. Akad. Wetensch. Indag. Math., 45(3):343–359, 1983. ISSN 0019-3577, 0023-3358.

Steinig:1972:SLF

[Ste72] J. Steinig. The sign of Lommel’s function. Transactions of the American Mathematical Society, 163:123–129, 1972. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Spigler:1990:AZS

[SV90a] R. Spigler and M. Vianello. Approximating zeros of solutions of second-order linear ODEs by “phase function” methods. In Wong [Won90], pages 707–722. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Spigler:1990:NME

[SV90b] Renato Spigler and Marco Vianello. A numerical method for evaluating zeros of solutions of second-order linear differential equations. Mathematics of Computation, 55(192):591–612, October 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spigler:1994:DCL

[SV94] Renato Spigler and Marco Vianello. Discrete and continuous Liouville–Green–Olver approximations: a unified treatment via Volterra–Stieltjes integral equations. SIAM Journal on Mathematical Analysis, 25(2):720–732, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/
article/23121.

Shivakumar:1988:EBU

[SW88] P. N. Shivakumar and R. Wong.

Error bounds for a uniform asymptotic expansion of the Legendre function P_{n}^{−m}(coshz).

Quarterly of Applied Mathematics, 46(3):473–488, 1988. CO-DEN QAMAAY. ISSN 0033-569X (print), 1552-4485 (electronic).

Swartzlander:1990:CAb

[Swa90] Earl E. Swartzlander, Jr. Computer Arithmetic, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8945-5. ix + 396 pp. LCCN QA76.9 .C62C66 1990. This is part of a twovolume collection of influential papers on the design of computer arithmetic.

Temme:1978:UAE

[Tem78] N. M. Temme. Uniform asymptotic expansions of confluent hypergeometric functions. Journal of the Institute of Mathematics and its Applications, 22(2):215– 223, 1978. CODEN JMTAA8. ISSN 0020-2932.

Temme:1982:UAE

[Tem82] N. M. Temme. The uniform asymptotic expansion of a class of integrals related to cumulative distribution functions. SIAM Journal on Mathematical Analysis, 13(2):239–253, March 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temme:1983:UAE

[Tem83] Nico M. Temme. Uniform asymptotic expansions of Laplace integrals. Analysis, 3(1–4):221–249, 1983. ISSN 0174-4747 (print), 2196-6753 (electronic).

Temme:1989:CSF

[Tem89] N. M. Temme. On the computation of special functions by using asymptotic expansions. In Numerical and applied mathematics, Part II (Paris, 1988), volume 1.2 of IMACS Ann. Comput. Appl. Math., pages 463–466. Baltzer, Basel, Switzerland, 1989. ISBN 3-905135-60-4.

Temme:1990:PAE

[Tem90] Nico M. Temme. Polynomial asymptotic estimates of Gegenbauer, Laguerre, and Jacobi polynomials. In Wong [Won90], pages 455–476. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Temme:1992:AII

[Tem92] N. M. Temme. Asymptotic inversion of incomplete gamma functions. Mathematics of Computation, 58(198):755–764, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thompson:2011:NHM

[Tho11] Ian Thompson. NIST Handbook of Mathematical Functions, edited by Frank W. J. Olver, Daniel W. Lozier, Ronald F. Boisvert, Charles W. Clark. Contemporary Physics, 52(5):497– 498, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Tremblay:1976:SIP

[TL76] Richard Tremblay and Marie-Louis Lavertu. Sur une série impliquant les puissances des zéros d’un polynôme. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences. Séries A et B, 282(3):Aii, A155–A157, 1976. CODEN CHASAP. ISSN 0151-0509.

Turner:1989:NAP

[Tur89] Peter R. Turner, editor. Numerical Analysis and Parallel Processing: Lectures given at The Lancaster Numerical Analysis Summer School 1987, volume 1397 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51645-X. vi + 264 pp. Lectures from the Third S.E.R.C. Numerical Analysis Summer School held at the University of Lancaster, Lancaster, 1987.

Temme:2014:Pa

[TW14a] Nico Temme and Roderick Wong. Preface. Analysis and Applications (Singapore), 12(4):vii–viii, June 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part I).

Temme:2014:Pb

[TW14b] Nico Temme and Roderick Wong. Preface. Analysis and Applications (Singapore), 12(5):vii, August 2014. ISSN 0219-5305 (print), 1793-6861 (electronic).

Special Issue: Dedicated to the Memory of Frank Olver (Part II).

Temme:2014:Pc

[TW14c] Nico Temme and Roderick Wong.

Preface. Analysis and Applications (Singapore), 12(6):vii–viii, October 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part III).

Ursell:1959:ORC

[UO61] F. Ursell and F. W. J. Olver.

Obituary: Roger Chapman Thorne. Journal of the Australian Mathematical Society, 1(2):255– 256, February 1959/61. CO-DEN JAUMAX. ISSN 0004-9735 (print), 2059-9234 (electronic).

Ursell:1987:UAE

[Urs87] F. Ursell. Uniformly asymptotic expansions for an integral with a large and a small parameter. Mathematical Proceedings of the Cambridge Philosophical Society, 101(2):349–362, 1987. CO-DEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Ursell:1990:ILP

[Urs90]  F. Ursell.  Integrals with

a large parameter and the maximum-modulus principle. In Wong [Won90], pages 477–489. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

VanderCruyssen:1979:ROA

[Van79] P. Van der Cruyssen. A reformulation of Olver’s algorithm for the numerical solution of secondorder linear difference equations. Numerische Mathematik, 32(2): 159–166, June 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verosky:1991:NPO

[Ver91] John M. Verosky. Negative powers of Olver recursion operators. Journal of Mathematical Physics, 32(7):1733–1736, July 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

                                        aip.org/resource/1/jmapaq/
                                        v32/i7/p1733_s1.

Vrahatis:1996:LCZ

[VRZG96] M. N. Vrahatis, O. Ragos, F. A. Zafiropoulos, and T. N. Grapsa. Locating and computing zeros of Airy functions. Zeitschrift für Angewandte Mathematik und Mechanik, 76(7):419–422, 1996. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Walker:1991:IDG

[Wal91] Peter Walker. Infinitely differentiable generalized logarithmic and exponential functions. Mathematics of Computation, 57(196): 723–733, October 1991. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wan:1990:FAE

[Wan90] Frederic Y. M. Wan. Finite axial extension and torsion of elastic helicoidal shells. In Wong [Won90], pages 491–516. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Wazwaz:2007:NSK

[Waz07]  Abdul-Majid Wazwaz.  New

solitons and kinks solutions to the Sharma–Tasso–Olver equation. Applied Mathematics and Computation, 188(2):1205– 1213, May 15, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/S0096300306014640.

Wazwaz:1986:MUA

[WH86a] Abdul-Majid Wazwaz and Floyd B.

Hanson. Matched uniform approximations for a singular boundary point and an interior turning point. SIAM Journal on Applied Mathematics, 46(6): 943–961, December 1986. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712x (electronic).

Wazwaz:1986:SBR

[WH86b] Abdul-Majid Wazwaz and Floyd B.

Hanson. Singular boundary resonance with turning point resonance. SIAM Journal on Applied Mathematics, 46(6):962–977, December 1986. CODEN SMJMAP.

ISSN 0036-1399 (print), 1095-712x (electronic).

Wilcox:1964:ASD

[Wil64] Calvin H. Wilcox, editor. Asymptotic solutions of differential equations and their applications: Proceedings of a Symposium Conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, May 4–6, 1964. Wiley, New York, NY, USA, 1964. Publication No. 13 of the Mathematics Research Center, United States Army, The University of Wisconsin. Dedicated to Professor Rudolph E. Langer.

Wimp:1975:BRA

[Wim75] Jet Wimp. Book review: Asymptotics and Special Functions (F. W. J. Olver). SIAM Review, 17 (3):569–575, ???? 1975. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Wimp:1980:USF

[Wim80] Jet Wimp. Uniform scale functions and the asymptotic expansion of integrals. Lecture Notes in Mathematics, 827: 251–271, 1980. CODEN LN-MAA2. ISBN 3-540-10252-3 (print), 3-540-38346-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

                                        /link.springer.com/chapter/
                                        10.1007/BFb0091383/.

Wimp:1990:SPC

[Wim90] Jet Wimp. Some properties of convolution sequences and asymptotics for the Taylor coefficients for products of Bessel functions. In Wong [Won90], pages 517–533. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Wong:1983:ASR

[Won83] R. Wong. Applications of some recent results in asymptotic expansions. Congressus Numerantium, 37:145–182, 1983. ISSN 0384-9864.

Wong:1989:SRR

[Won89] R. Wong. Some recent results in uniform asymptotic expansions of special functions. In Numerical and applied mathematics, Part II (Paris, 1988), volume 1.2 of IMACS Ann. Comput. Appl. Math., pages 473–477. Baltzer, Basel, Switzerland, 1989. ISBN 3-905135-60-4.

Wong:1990:ACA

[Won90] R. (Roderick) Wong, editor.

Asymptotic and Computational Analysis: Conference in Honor of Frank W. J. Olver’s 65th Birthday, volume 124 of Lecture Notes in Pure and Applied Mathematics. Marcel Dekker, New York, NY, USA, 1990. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Wong:2000:SPF

[Won00] Roderick Wong, editor. Selected papers of F. W. J. Olver. Part I, II, volume 7 of World Scientific Series in 20th Century Mathematics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2000. ISBN 981-02-4106-2. xviii + 520 (Part I); x + 521– 1074 (Part II) pp. URL https:

                                        //www.worldscientific.com/
                                        worldscibooks/10.1142/4251.

Wong:2013:OFW

[Won13]  Roderick Wong.  Obituaries:

Frank W. J. Olver, 1924– 2013. SIAM News Web site,

                                        2013. URL https://sinews.
                                        siam.org/About-the-Author/
                                        obituaries-frank-wj-olver-
                                        1924-2013.

Wong:2014:ALR

[Won14a] R. Wong. Asymptotics of linear recurrences. Analysis and Applications (Singapore), 12(4):463– 484, June 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part I).

Wong:2014:MFW

[Won14b] R. Wong. In memoriam Frank W. J. Olver (1924–2013). Analysis and Applications (Singapore), 12(4):xi–xxvi, June 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part I).

Wong:1980:UAE

[Won81] R. Wong. On a uniform asymptotic expansion of a Fouriertype integral. Quarterly of Applied Mathematics, 38(2):225– 234, 1980/81. CODEN QA-MAAY. ISSN 0033-569X (print), 1552-4485 (electronic).

Wood:1990:EES

[WP90] A. D. Wood and R. B. Paris.

On eigenvalues with exponentially small imaginary part. In Wong [Won90], pages 741–749. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Wang:2023:GCC

[WYX^{+}23] Ziyu Wang, Haobin Yang, Danlin Xu, Zehong Liang, Lingling Shui, and Dongmei Deng. Generation and control of the circle Olver beams. Optics Express, 31 (4):6241, February 2023. CO-DEN OPEXFF. ISSN 1094-4087.

Young:1960:BFP

[YK60] A. Young and A. Kirk, editors. Bessel functions: (Part 4: Kelvin functions), volume 7, 10 of Royal Society Mathematical Tables. Cambridge University Press, Cambridge, UK, 1960. ???? pp.

Yao:2012:FST

[YL12] Biwei Yao and Senyue Lou.

Fermionization of Sharma–Tasso– Olver system. Chinese Annals of Mathematics, Series B, 33(2): 271–280, March 2012.

Yue:2014:AGS

[YX14] Chao Yue and Tiecheng Xia. Algebro-geometric solutions for the complex Sharma–Tasso– Olver hierarchy. Journal of Mathematical Physics, 55(8):083511, August 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zayed:2016:MSE

[ZAAN16] Elsayed M. E. Zayed, Yaser A. Amer, and Abdul-Ghani Al-Nowehy. The modified simple equation method and the multiple exp-function method for solving nonlinear fractional Sharma– Tasso–Olver equation. Acta Mathematicae Applicatae Sinica, English Series, 32(4):793–812, October 2016. ISSN 0168-9673 (print), 1618-3932 (electronic).

Zahar:1990:GOA

[Zah90] R. V. M. Zahar. A generalization of Olver’s algorithm for linear difference systems. In Wong [Won90], pages 535–551. ISBN 0-8247-8347-6. LCCN QA299.6 .A88 1990. URL

http://www.loc.gov/catdir/
enhancements/fy0647/90002810-
d.html.

Zanovello:1995:NAS

[Zan95] Renato Zanovello. Numerical analysis of Struve functions with applications to other special functions. Annals of Numerical Mathematics, 2(1–4):199–208, 1995. ISSN 1021-2655. Special functions (Torino, 1993).

Zayed:2011:NMS

[Zay11] Elsayed M. E. Zayed. A note on the modified simple equation method applied to Sharma– Tasso–Olver equation. Applied Mathematics and Computation, 218(7):3962–3964, December 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/S009630031101191X.